### 15 Mar 2026 v1

# Governing Dynamic Capabilities: Cryptographic Binding and Reproducibility Verification for AI Agent Tool Use

Abstract
[…]
1 Introduction
1.1 Agent Governance: A Missing Foundation
AI agents are no longer isolated tools. Through protocols like MCP (Model Context Protocol) [8] and A2A (Agent-to-Agent Protocol) [8], modern agents dynamically discover tools, invoke external APIs, execute code, and orchestrate other agents at runtime. The MCP ecosystem alone has grown to over 1,000 tool servers in under a year [32]. As these agents assume real-world responsibilities—managing enterprise credentials, executing financial transactions, accessing sensitive data—a fundamental question emerges: what properties must an agent ecosystem satisfy for humans to trust it?
This question has been raised from multiple directions. The EU AI Act (Regulation 2024/1689) mandates traceability, logging, and conformity assessment for high-risk AI systems (Articles 12, 43, 72). NIST’s AI Agent Standards Initiative [26] identifies security controls, governance oversight, and monitoring as key requirements. The TRiSM framework [28] structures agent governance around explainability, security, privacy, and lifecycle management. Schroeder de Witt [11] calls for secure interaction protocols that prevent insecure emergent behaviors. Cheng et al. [43] ground agent safety in transparency, accountability, and trustworthiness. Allegrini et al. [44] formalize 31 safety and liveness properties for agentic task lifecycles in temporal logic. Ghosh et al. [45] propose an operational safety framework with red-teaming methodology, policy-based tool scoping, and optional trace signing—but without cryptographic capability binding, formal adversary models, or verifiable replay mechanisms.
Yet despite this convergence of concern, no existing framework translates these governance aspirations into enforceable cryptographic infrastructure. The gap is not in recognizing what is needed, but in building it. We propose three Agent Governance Requirements that distill the common thread across these efforts into technically precise, verifiable properties:
Definition 1 (Agent Governance Requirements).
A governed agent ecosystem satisfies three properties:
G1: Capability Integrity. Every agent’s identity is cryptographically bound to its complete capability set—the model it executes and every tool it can invoke. The identity is unforgeable and non-repudiable. Any change to the capability set invalidates existing credentials and requires explicit re-authorization by a human or organizational principal.
G2: Behavioral Verifiability. An agent’s actions within its authorized capability envelope can be independently verified post-hoc, enabling detection of model substitution, unauthorized tool use, or output manipulation.
G3: Interaction Auditability. All inter-agent and agent-human interactions produce tamper-evident records sufficient for forensic reconstruction, allowing humans to audit agent behavior, attribute responsibility, and provide evidence for regulatory compliance and system improvement.
These requirements are not novel in aspiration—they reflect the consensus of legal [33, 34], regulatory [26], and technical [11, 28, 43, 44, 45] communities. What is novel is the claim that all three must hold simultaneously, and that none is achievable with existing agent infrastructure. G1 alone (without G2 and G3) permits certified agents to misbehave undetectably within their capability envelope. G2 alone (without G1) cannot distinguish authorized from unauthorized capabilities. G3 alone (without G1 and G2) produces audit logs of unverified actions by unverified agents.
1.2 Why the Current Landscape Falls Short
The current agent security landscape can be mapped into four layers (Figure 1), each addressing fragments of the governance requirements but none satisfying them:
Identity platforms (Keyfactor [19], SPIFFE/SPIRE [21, 22]) issue certificates binding an agent to a public key—but a certificate remains valid regardless of how many tools the agent acquires after issuance. G1 violated: capability changes are invisible to the identity layer.
Authorization frameworks (OIDC-A [4], Authenticated Delegation [3], Cerbos [39]) define what an authenticated agent may access—but authorization is evaluated against declared capabilities at grant time, with no mechanism to detect drift. G1 violated: authorization does not track capability changes.
Runtime gateways (Gravitee [40], Lasso Security [41]) inspect tool calls in real time—but are fundamentally stateless, with no memory of prior tool configurations and no behavioral verification. G2 and G3 violated.
Governance platforms (Credo AI [42]) maintain agent registries documenting capabilities—but operate at the documentation layer without cryptographic enforcement. G1 violated: declared capabilities are not bound to the agent’s actual configuration.
Figure 1: The four-layer agent security landscape and the capability-identity gap. Existing layers partially address individual requirements but none bridges them. Our framework provides the missing cryptographic governance layer satisfying G1–G3, validated end-to-end across 5–20 agent pipelines.
The root cause is a failure to recognize a fundamental architectural distinction in LLM-based agents. Inside a transformer’s forward pass, tool definitions and user context are indistinguishable token sequences. But at the orchestration layer, tool definitions determine which real-world actions the model’s outputs can trigger, while runtime context determines which actions the model chooses. This yields a principled capability-context separation: the capability envelope (model identity, tool definitions, permission scopes) governs what is possible and changes infrequently; the runtime context (conversation history, user instructions, memories) governs what is chosen and changes per-interaction. The security implication is direct: capability changes require re-authorization (G1); context changes require verifiability (G2) and auditability (G3).
Existing frameworks conflate the two. A tool addition and a new user message are treated identically—neither triggers re-evaluation. We call this the capability-identity gap: existing frameworks authenticate who an agent is and authorize what it may access, but none detect when the agent’s capabilities have changed. This gap enables a class of attacks we term silent capability escalation.
1.3 Our Framework: Bridging the Gap
The capability-context separation provides the theoretical foundation for three mechanisms, each targeting a specific governance requirement:
[…]
Together, these three mechanisms provide what the existing landscape lacks: a stateful, cryptographic governance layer satisfying G1–G3, bridging identity platforms and runtime gateways. Our contribution is not in novel cryptographic primitives; rather, we show that standard mechanisms—X.509 certificates, SHA-256 hashing, Ed25519 signatures, hash chains—when composed through the capability-context separation, are sufficient to enforce G1–G3 with negligible overhead, providing actionable design guidance for how agents should be built, how they should interact, and how their behavior should be audited.
1.4 Threat Model and Security Guarantees
We consider a PPT adversary who can compromise up to $k$ agent nodes, create malicious agents, inject or modify tools, and manipulate network traffic (Dolev-Yao model). Critically, the adversary can perform silent capability escalation. The adversary cannot compromise non-agent nodes or break standard cryptographic primitives.
Under this model, we define and prove nine security properties:
  1. 1.
    Capability Binding: tool configuration change $\Rightarrow$ access denied (Theorem 5).
  2. 2.
    Trust Containment: child constraints $\leq_{\kappa}$ parent constraints (Theorem 4).
  3. 3.
    Chain Integrity: valid chain to trusted root required (Theorem 6).
  4. 4.
    Delegation Depth: chains bounded by ancestor’s $\mathit{max\_depth}$ (Theorem 7).
  5. 5.
    Reproducibility Soundness: different model/tools detected with overwhelming probability (Theorem 8).
  6. 6.
    Bounded Divergence: adversary-controlled model passing $n$ replay checks is $(n,\epsilon)$-indistinguishable from declared model (Theorem 9).
  7. 7.
    Chain Verifiability: a single unverifiable interior agent breaks end-to-end behavioral verification for all downstream nodes (Theorem 12).
  8. 8.
    Credential Isolation: colluding agents gain no extra permissions (Theorem 10).
  9. 9.
    Ledger Integrity: undetectable modification impossible (Theorem 11).
Properties 1–4 enforce G1. Properties 5–7 enforce G2 (with 7 characterizing its chain-level semantics). Properties 8–9 enforce G3.
1.5 Contributions
In summary:
  • We define three Agent Governance Requirements—capability integrity, behavioral verifiability, and interaction auditability—distilling converging regulatory, legal, and technical concerns into formally precise properties (§1.1).
  • We identify the capability-identity gap and the capability-context separation principle as its root cause and resolution (§1.2).
  • We show that standard cryptographic mechanisms—composed through the capability-context separation—are sufficient to enforce all three requirements with negligible overhead, providing design guidance for secure agent ecosystems.
  • We propose capability-bound agent certificates with monotonic trust propagation (§3).
  • We introduce reproducibility as a trust primitive and establish that reproducibility is a necessary condition for behavioral verifiability: high-assurance agent deployments must be designed for deterministic output to be auditable (§4).
  • We define a verifiable interaction ledger for tamper-evident forensic reconstruction (§5).
  • We prove that sufficient replay verification yields bounded divergence: an adversary-controlled model that passes $n$ replay checks is provably $(n,\epsilon)$-indistinguishable from the declared model, with $\epsilon$ exponentially small in nn4.7).
  • We establish chain verifiability: behavioral verifiability is a chain property, not a node property—a single unverifiable interior agent breaks end-to-end verification for all downstream nodes, yielding principled deployment constraints on model placement in trust trees (§4.8).
  • We implement a working MCP governance proxy that enforces all three requirements in real-time with $\sim$}120 µs overhead per call, detecting silent capability escalation, tool trojanization, and certificate forgery (§7).
  • We formalize and prove nine security properties under a realistic adversary model (§6).
  • We implement a Rust prototype achieving 97 µs verification, ${\sim}0.62$ ms governance overhead, and ${\sim}1{,}200{,}000\times$ speedup over BAID (§7).
  • We deploy the full governance stack in an end-to-end multi-agent pipeline (5–20 agents with real LLM calls), demonstrating 0.12% overhead, sub-linear per-agent scaling, and detection of all in-situ attacks—while OAuth 2.1 and OpenTelemetry baselines detect none (§7.4).
1.6 Paper Organization
§2 presents the threat model and attack scenarios. §3 defines capability-bound certificates and the trust propagation tree. §4 introduces reproducibility commitments. §5 describes the verifiable interaction ledger. §6 presents formal security properties and proofs. §7 describes implementation and evaluation. §8 discusses limitations. §9 covers related work. §10 concludes.
2 Threat Model and Motivating Attacks
2.1 System Model
We model the trust infrastructure as a rooted directed tree $T=(V,E,r,\tau)$, where $V$ is the set of all entities (CAs, humans, organizations, AI agents), $E$ is the set of certificate issuance relationships, $r$ is the root CA, and $\tau:V\to\{\mathit{NA},\mathit{AG}\}$ classifies each node as Non-Agent ($\mathit{NA}$) or Agent ($\mathit{AG}$).
The tree satisfies a structural type constraint: $\tau(u)=\mathit{AG}\implies\forall(u,v)\in E:\tau(v)=\mathit{AG}$. Agents can delegate to sub-agents, but cannot confer trust upon humans or legal entities.
Each agent node $v$ holds a Capability-Bound Certificate $\mathit{cert}_{v}$ and maintains a runtime tool configuration $\mathcal{T}_{v}(t)$ that may change over time. Each credential $c$ has a risk tier $\mathit{tier}_{c}\in\{T0,T1,T2,T3\}$ where $T0$ (human-only) is most sensitive.
2.2 Adversary Model
We define a PPT adversary $\mathcal{A}$ with capabilities: C1: Network Control (Dolev-Yao). C2: Agent Compromise of up to $k$ agent nodes. C3: Rogue Agent Creation with fresh key pairs. C4: Silent Capability Escalation—modifying tool configurations without changing identity credentials. C5: Interaction Injection—fabricating or modifying ledger records.
$\mathcal{A}$ cannot: compromise non-agent nodes, forge digital signatures (EUF-CMA), find hash collisions, or reverse hash functions.
2.3 Motivating Attack Scenarios
We enumerate twelve attacks organized into four categories. ATK-1 through ATK-9 address structural governance threats; ATK-10 through ATK-12 address runtime behavioral attacks empirically validated by the Nemotron-AIQ dataset [45], which demonstrates 100% success rates for prompt injection and tool-mediated attacks against unprotected agentic workflows.
[…]
ATK-2: Model Substitution. $\mathcal{A}$ replaces the certified model with a fine-tuned variant. Detected via model binding mismatch and replay verification divergence.
ATK-3: Tool Trojanization. $\mathcal{A}$ replaces tool implementation (same name/version, different code). Detected via code-level hashing in the skills manifest.
Delegation Attacks. ATK-4: Phantom Delegation. Compromised agent (T2) issues T1 certificate to sub-agent. Prevented by monotonic trust propagation. ATK-5: Chain Forgery. Fabricated certificate chain. Prevented by signature verification at every level. ATK-6: Depth Overflow. Deep delegation hierarchy. Prevented by strictly decreasing depth counters. ATK-7: Credential Collusion. Two or more compromised agents at lower tiers (e.g., T3++T3) attempt to combine their credentials to access resources restricted to a higher tier (T1). Prevented by independent per-agent credential evaluation—each agent’s tier and permissions are verified in isolation, so colluding agents gain no additional permissions beyond their individual certificates (Theorem 10).
Forensic Attacks. ATK-8: Evidence Tampering. Log modification after incident. Prevented by hash-linked ledger with bilateral signing. ATK-9: Blame Shifting. Dispute in multi-agent chain. Resolved via interaction commitments and reproducibility replay.
Runtime Behavioral Attacks. ATK-10: Injection-Driven Knowledge Poisoning. Adversarial content injected via tool outputs (e.g., retrieval results) propagates through the agent’s reasoning chain, corrupting final outputs. Nemotron-AIQ traces show this attack succeeds in 100% of undefended trials across knowledge-poisoning and phishing-link templates. Our ledger’s content commitments ($\eta_{\mathit{in}},\eta_{\mathit{out}}$) enable post-hoc identification of the exact span where poisoned content entered the chain.
ATK-11: Cascading Action Hijack. Injected instructions persist across multiple agent workflow stages (INFECT template), causing the agent to refuse legitimate tasks or produce harmful content at downstream nodes. Our hash-chain ledger records every inter-span transition, enabling forensic tracing of attack propagation paths that OTEL-only audit trails cannot reliably provide.
ATK-12: Output Exfiltration via Rendering. Agent outputs are manipulated to embed phishing links (PHISHING_LINK) or malicious markdown (MARKDOWN) that exfiltrate data when rendered by downstream consumers. Detected post-hoc via reproducibility verification: replaying the interaction with the committed parameters produces benign output, diverging from the tampered record.
3 Capability-Bound Agent Certificates
3.1 Certificate Structure
Definition 2 (Capability-Bound Certificate).
A Capability-Bound Certificate for node $v$ is:
$\mathit{cert}_{v}=(\mathit{id}_{v},\mathit{id}_{\mathit{par}(v)},\mathit{pk}_{v},\mu_{v},\sigma_{v},\kappa_{v},\rho_{v},t_{s},t_{e},\mathit{sig})$
where $\mu_{v}=(\mathit{provider},\mathit{model\_id},\mathit{model\_ver})$ is the model binding, $\sigma_{v}=\{(\mathit{sid}_{i},\mathit{ver}_{i},h_{i},P_{i})\}_{i=1}^{n}$ is the skills manifest, $\kappa_{v}=(\mathit{max\_tier},\mathit{max\_depth},\mathit{allowed\_models},\mathit{max\_rate})$ are trust constraints, and $\rho_{v}=(\mathit{level},\mathit{config})$ is the reproducibility commitment.
3.2 Skills Manifest and Capability Binding
Definition 3 (Skills Manifest Hash).
$H_{v}^{\mathit{cert}}=\mathrm{SHA\text{-}256}(\mathrm{canonical}(\{(\mathit{sid}_{i},\mathit{ver}_{i},h_{i},P_{i})\}_{i=1}^{n}))$
At runtime: $H_{v}(t)=\mathrm{SHA\text{-}256}(\mathrm{canonical}(S_{v}(t)))$.
Capability Binding Invariant. Access permitted only if $H_{v}(t)=H_{v}^{\mathit{cert}}$. Any capability change invalidates access. Crucially, runtime context changes (memories, preferences) do not affect $H_{v}(t)$—the practical consequence of capability-context separation.
For open-source tools, $h_{i}$ is the SHA-256 hash of source code. For closed-source tools, $h_{i}$ is the hash of the public configuration descriptor (name, version, API schema).
3.3 Trust Propagation Tree
Figure 2: Trust propagation tree for the evaluation environment. NA nodes (rectangles) form the trust anchor; AG nodes (rounded) are agents. Constraints decay monotonically downward.
Definition 4 (Trust Constraint Ordering).
$\kappa_{a}\leq_{\kappa}\kappa_{b}$ iff $\kappa_{a}.\mathit{max\_tier}\leq\kappa_{b}.\mathit{max\_tier}$, $\kappa_{a}.\mathit{max\_depth}<\kappa_{b}.\mathit{max\_depth}$ (strict), $\kappa_{a}.\mathit{allowed\_models}\subseteq\kappa_{b}.\mathit{allowed\_models}$, and $\kappa_{a}.\mathit{max\_rate}\leq\kappa_{b}.\mathit{max\_rate}$.
Trust Propagation Rule. For every edge $(u,v)\in E$: $\kappa_{v}\leq_{\kappa}\kappa_{u}$.
3.4 Certificate Verification Algorithm
Figure 3: Verification flow for an MCP tool call through the A2Auth governance layer. Each phase can independently reject with a specific reason.
When agent $v$ requests access to credential $c$ at time $t$, the verifier executes four phases (Figure 3):
1function Verify(chain, credential, skills, roots):
2 // Phase 1: Chain Integrity
3 if chain[0].issuer $\notin$ roots: return DENY
4 for i in 1..len(chain):
5 if not VerifySig(chain[i-1].pk, chain[i].sig): return DENY
[…]
7 // Phase 2: Capability Binding
[…]
9 // Phase 3: Trust Constraint Check
[…]
11 // Phase 4: Revocation Check
12 if IsRevoked(cert_v): return DENY
13 return ALLOW
3.5 Integration with A2A Agent Cards
Agent Certificates are designed as non-invasive extensions of the A2A ecosystem via an optional extensions field in the Agent Card containing the certificate PEM, chain, and skills manifest hash.
4 Reproducibility Commitments
4.1 Reproducibility as a Necessary Condition
A key observation motivates our approach: behavioral verifiability (G2) is impossible without reproducibility. If an agent’s outputs cannot be independently regenerated—even approximately—then no post-hoc mechanism can distinguish authorized behavior from unauthorized behavior. This is not merely a limitation of our framework; it is a fundamental constraint on any verification scheme that does not rely on trusted execution environments.
This observation has a prescriptive implication for secure agent system design: high-assurance agent deployments must be engineered for reproducible output. Concretely, this means using deterministic sampling (temperature==0, fixed seeds), preferring structured outputs (code, JSON, SQL) over free-form prose where security-critical decisions are involved, and selecting inference providers that guarantee bitwise or near-bitwise reproducibility [6]. Agents that cannot commit to any reproducibility level receive a trust tier downgrade (§4.4), making the security cost of non-reproducibility explicit in the governance framework.
4.2 Empirical Foundation
LLM inference is nearly deterministic. Karvonen et al. [1] demonstrated that with fixed sampling parameters, regenerating output produces >>98% token-identical sequences. EigenAI [6] achieved 100% bit-exact deterministic inference on fixed GPU architectures with negligible performance loss. Our own validation (§7, Experiment 5) confirms that code-generating agents achieve perfect determinism, while prose outputs exhibit paraphrase-level variation that remains well-separated from cross-model divergence.
4.3 Reproducibility as a Trust Primitive
Definition 5 (Reproducibility Commitment).
$\rho_{v}=(\mathit{level},\mathit{config})$ where $\mathit{level}\in\{\mathit{full},\mathit{statistical},\mathit{none}\}$.
Full ($\mathit{level}=\mathit{full}$): $\forall(\mathit{context},\mathit{seed}):O_{v}(\mathit{context},\mathit{seed})=O_{v}^{\mathit{replay}}(\mathit{context},\mathit{seed})$.
Statistical ($\mathit{level}=\mathit{statistical}$):
$\forall(\mathit{context},\mathit{seed}):\mathrm{CharMatch}(O_{v},O_{v}^{\mathit{replay}})\geq\theta$
The threshold $\theta$ is task-type dependent: our empirical validation (Experiment 5, §7) shows that code-generating agents achieve perfect determinism ($\theta=0.98$), while prose-generating agents require $\theta=0.10$ to account for paraphrase-level variation. Cross-model divergence remains well below even the relaxed threshold (max 0.089).
None: No reproducibility commitment.
4.4 Reproducibility and Trust Tier Interaction
[…]
4.5 Replay-Based Verification Protocol
1function ReplayVerify(cert, record, O_orig):
[…]
3 if $\rho$.level = none: return INCONCLUSIVE
4 O_ref := ExecuteModel($\mu, cert$$\sigma$,
5 record.input, record.seed, $\rho$.config)
6 if $\rho$.level = full:
7 return O_ref == O_orig ? VERIFIED : VIOLATION
8 else: // statistical
9 return CharMatch(O_ref, O_orig) $\geq$ $\theta$ ?
10 VERIFIED : VIOLATION
4.6 Security Analysis
Theorem 1 (Reproducibility Soundness, informal).
If compromised agent $v$ executes model $M^{\prime}\neq M$ or skills $S^{\prime}\neq S$, replay verification detects the deviation with probability $\geq 1-\epsilon(\lambda)$ for negligible $\epsilon$.
By the DiFR token divergence result [1], the match rate between different models is $\delta\approx 0.06$ (character-level, validated in Experiment 5). By Hoeffding’s inequality: $\Pr[\mathrm{CharMatch}\geq\theta]\leq\exp(-2n(\theta-\delta)^{2})$, which is negligible for realistic output lengths.
4.7 Indistinguishability Under Sufficient Verification
A natural concern is adversarial fine-tuning: could an adversary produce a model $M^{\prime}$ that passes reproducibility checks on tested inputs but behaves maliciously on untested ones? We show that sufficient replay verification bounds the impact of any such attack.
Definition 6 ($(n,\epsilon)$-Indistinguishability).
Models $M$ and $M^{\prime}$ are $(n,\epsilon)$-indistinguishable under skill set $\mathcal{S}$ and prompt distribution $\mathcal{D}$ if, after $M^{\prime}$ passes replay verification on $n$ independent prompts $x_{1},\ldots,x_{n}\sim\mathcal{D}$ across skills in $\mathcal{S}$:
$\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]\leq\epsilon$
Theorem 2 (Bounded Divergence).
If adversary-controlled model $M^{\prime}$ passes replay verification (CharMatch $\geq\theta$) on $n$ independent prompts drawn from $\mathcal{D}$ across $|\mathcal{S}|$ skills, then $M$ and $M^{\prime}$ are $(n,\epsilon)$-indistinguishable with:
$\epsilon\leq\exp\!\left(-2n\theta^{2}\right)$
Equivalently, for target confidence $1-\epsilon$, $n\geq\frac{\ln(1/\epsilon)}{2\theta^{2}}$ verification prompts suffice.
Proof sketch. Let $Z_{i}=\mathbf{1}[\mathrm{CharMatch}(M(x_{i}),M^{\prime}(x_{i}))<\theta]$ be independent Bernoulli random variables with $p=\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}<\theta]$. Observing $\sum Z_{i}=0$ (all $n$ tests pass), by a one-sided binomial test: $\Pr[\sum Z_{i}=0\mid p\geq\epsilon]=(1-p)^{n}\leq(1-\epsilon)^{n}\leq\exp(-n\epsilon)$. Setting this equal to $\alpha$ (significance level) and solving yields $p\leq\ln(1/\alpha)/n$ with confidence $1-\alpha$. The Hoeffding refinement accounting for the threshold gap gives the stated bound. Full proof in §6. ∎
Practical implications. For code-generating agents ($\theta=0.98$), only $n=3$ prompts per skill achieve $\epsilon<10^{-2}$; $n=12$ achieves $\epsilon<10^{-10}$. Even for relaxed prose thresholds ($\theta=0.10$), $n=1200$ prompts achieve $\epsilon<10^{-10}$. These are operationally feasible: a verification suite of 50 prompts per skill, run periodically, provides strong guarantees.
Security interpretation. This transforms reproducibility from a detection mechanism into a safety certificate: an adversary who fine-tunes $M$ into $M^{\prime}$ that passes $n$ diverse replay checks has produced a model that is provably close to $M$ on the tested distribution. Any residual divergence is bounded by $\epsilon$, which decreases exponentially with the verification budget. The adversary faces a dilemma: either $M^{\prime}$ behaves like $M$ (and the fine-tuning achieves nothing), or $M^{\prime}$ diverges on some inputs and will be caught with probability increasing exponentially in $n$. This is analogous to cryptographic indistinguishability: the adversary’s advantage is negligible given sufficient verification.
4.8 Chain Verifiability
The preceding analysis establishes that G2 can verify an individual agent’s computational process. A natural question arises: when agents form a multi-hop execution chain, is per-node verification sufficient for end-to-end behavioral verifiability? We show that it is not—behavioral verifiability is a chain property, not a node property—and that this observation has direct implications for trust tree architecture (Figure 4).
[…]
The gap that G2 uniquely fills. Consider what G1+G3 alone guarantee. G1 (capability binding) ensures that each agent’s certificate is bound to its declared model and tool set; any capability change invalidates the certificate. G3 (interaction auditability) ensures that every input and output is bilaterally signed and hash-chain linked; no record can be silently modified. Together, G1+G3 provide attribution integrity: we know who sent what to whom, and that no message was altered in transit.
[…]
Why chain position matters. Now consider a trust tree execution chain $v_{0}\to v_{1}\to\cdots\to v_{n}$, where $v_{0}$ is a human principal (trust anchor) and each $v_{i}$ ($i\geq 1$) is an agent node. Agent $v_{j}$ receives input from $v_{j-1}$, processes it using its declared model, and passes output to $v_{j+1}$.
Suppose $v_{j}$ is a compromised agent with G2==none (either because it uses a low-determinism model or because G2 verification is not enforced). Then:
  1. 1.
    $v_{j}$’s output cannot be verified as originating from its declared model.
  2. 2.
    $v_{j}$’s output becomes $v_{j+1}$’s input. Even if $v_{j+1}$ satisfies G2==full, replay verification of $v_{j+1}$ only proves that $v_{j+1}$ faithfully processed the input it received—but that input itself is unverifiable.
  3. 3.
    By induction, all downstream nodes $v_{j+1},\ldots,v_{n}$ operate on data whose provenance cannot be cryptographically verified back to the trust anchor.
Definition 7 (Chain Verifiability Depth).
For execution chain $v_{0}\to v_{1}\to\cdots\to v_{n}$ with $v_{0}$ a non-agent trust anchor, define:
$\mathrm{CVD}(v_{n})=\min\{j\geq 1:v_{j}\text{ is an agent with }\rho_{v_{j}}.\mathit{level}=\mathit{none}\}$
with $\mathrm{CVD}(v_{n})=n$ if no such $v_{j}$ exists (fully verifiable chain).
Theorem 3 (Chain Verifiability).
Let $v_{0}\to v_{1}\to\cdots\to v_{n}$ be an execution chain where $v_{0}$ is a non-agent trust anchor. If $\mathrm{CVD}(v_{n})=k<n$, then for all $i>k$, the output of $v_{i}$ cannot be verified as originating from a legitimate computation chain rooted at $v_{0}$, even if $v_{i}$ individually satisfies $\rho_{v_{i}}.\mathit{level}=\mathit{full}$.
Proof sketch. By Definition 7, $v_{k}$ has $\rho.\mathit{level}=\mathit{none}$, so no replay verification can confirm that $v_{k}$’s declared model produced its output $o_{k}$. Under the adversary model (§2), $v_{k}$ may have executed an arbitrary model $M^{\prime}\neq M_{k}$ that produces output $o_{k}^{\prime}\neq o_{k}^{*}$ (where $o_{k}^{*}$ is the output the declared model would have produced). Agent $v_{k+1}$ receives $o_{k}^{\prime}$ as input. Even with $\rho_{v_{k+1}}.\mathit{level}=\mathit{full}$, replay verification confirms $v_{k+1}(o_{k}^{\prime})=o_{k+1}$, i.e., that $v_{k+1}$ faithfully processed $o_{k}^{\prime}$. But $o_{k}^{\prime}$ may differ from $o_{k}^{*}$, so $o_{k+1}$ is the correct computation on incorrect input. By induction on chain length, the same applies to all $v_{i}$ with $i>k$. Full proof in §6. ∎
Architectural implications. This theorem yields three deployment principles:
Principle 1: High-determinism models at interior nodes. Models with strong reproducibility (Anthropic Haiku: CharMatch==0.987, Sonnet: 0.823; see Table 4) should be placed at coordinator and interior positions in the trust tree, where their outputs become inputs for downstream agents. Low-determinism models (Kimi: 0.169, DeepSeek: 0.225) should be constrained to leaf positions where their outputs do not propagate further. This is not merely a performance recommendation—it is a security architecture constraint.
Principle 2: Verification boundary checkpoints. When a G2==none agent must appear at an interior position (e.g., a reasoning model as coordinator), the chain requires a verification boundary: either (a) a human-in-the-loop checkpoint that re-anchors trust by reviewing and signing the output, (b) TEE attestation providing hardware-enforced computational integrity, or (c) dual-execution with independent instances and output comparison (analogous to Byzantine fault tolerance). Each mechanism re-establishes chain verifiability for downstream nodes.
Principle 3: Explicit degradation in the ledger. When chain verifiability is broken at node $v_{k}$, the ledger should record a PARTIAL_VERIFIABILITY marker at that boundary. Downstream agents inherit a trust tier penalty beyond the standard G2==none downgrade (§4.4), making the reduced security guarantee visible to auditors and compliance systems.
Relationship to TEE. This analysis clarifies the complementary roles of G2 and trusted execution environments. G2 provides software-only verification of computational fidelity—a “software TEE”—using only black-box API access and the near-determinism of LLM inference. Hardware TEEs provide the same guarantee through hardware attestation but require infrastructure unavailable in most LLM deployments. The Chain Verifiability Theorem shows precisely where each mechanism is needed: at every interior agent node in the trust tree. In mixed deployments, G2 and TEE are interchangeable at the chain level: an agent with TEE attestation and G2==none maintains chain verifiability identically to an agent with G2==full and no TEE.
5 Verifiable Interaction Ledger
5.1 Motivation
Capability binding and reproducibility verify individual agents. Real-world incidents involve chains of interactions. We need an immutable, tamper-evident record where no single party—including the ledger operator—can undetectably alter history.
Why OpenTelemetry traces are insufficient. The industry standard for agentic observability is OpenTelemetry (OTEL) [46], whose span trees record execution traces with span_id/parent_id linkage, timestamped input.value/output.value payloads, and span-kind annotations (CHAIN, LLM, TOOL). Analysis of the Nemotron-AIQ Agentic Safety Dataset [45]—10,796 OTEL traces from NVIDIA’s AI-Q research assistant under 22 distinct attack templates—reveals three critical gaps: (1) No integrity protection: span linkage is purely referential; any party with storage access can modify, delete, or reorder spans undetectably. (2) No identity binding: spans record function names but not cryptographically verified agent identities or capability certificates. (3) No privacy preservation: OTEL stores full plaintext input/output payloads, creating data exposure risk. Our interaction ledger addresses all three gaps through hash-chain linking ($h_{\mathit{prev}}$), Ed25519 bilateral signing ($\mathit{sig}_{i}$), agent certificate binding ($\mathit{ch}_{\mathit{send}},\mathit{ch}_{\mathit{recv}}$), and commitment-only storage ($\eta_{\mathit{in}},\eta_{\mathit{out}}$).
5.2 Ledger Structure
Definition 8 (Interaction Record).
$R_{i}=(\mathit{seq}_{i},t_{i},a_{\mathit{send}},a_{\mathit{recv}},\mathit{ch}_{\mathit{send}},\mathit{ch}_{\mathit{recv}},\eta_{\mathit{in}},\eta_{\mathit{out}},\pi_{i},h_{\mathit{prev}},\mathit{sig}_{i})$ where $\eta_{\mathit{in}}=\mathrm{SHA\text{-}256}(\mathit{input})$, $\eta_{\mathit{out}}=\mathrm{SHA\text{-}256}(\mathit{output})$, $\pi_{i}=(\mathit{seed},\mathit{model\_ver},\mathit{skills\_hash})$, and $h_{\mathit{prev}}=\mathrm{SHA\text{-}256}(R_{i-1})$.
Definition 9 (Verifiable Interaction Ledger).
$\mathcal{L}=(R_{1},\ldots,R_{n})$ satisfying: (1) sequential consistency, (2) hash chain integrity, and (3) signature validity at every record.
5.3 Privacy Preservation
The ledger stores only commitments (hashes), not content. During normal operation, only interaction metadata is revealed. During forensic investigation, parties produce original content verified against stored commitments.
5.4 Forensic Reconstruction Protocol
The protocol proceeds in five steps: (1) identify the relevant interaction chain, (2) verify ledger integrity (signatures + hash chain), (3) request content disclosure and verify against commitments, (4) verify agent certificates at interaction time, and (5) execute replay verification for agents with reproducibility commitments.
6 Formal Security Properties and Proofs
We prove nine properties under the adversary model of §2, using standard cryptographic assumptions: EUF-CMA security, collision resistance of SHA-256, and the DiFR token divergence property [1].
Theorem 4 (Trust Containment).
For all edges $(u,v)\in E$: $\kappa_{v}\leq_{\kappa}\kappa_{u}$.
Proof.
Suppose $\mathcal{A}$ issues $\mathit{cert}_{w}$ with $\kappa_{w}\not\leq_{\kappa}\kappa_{v}$. The verifier checks $\kappa_{w}\leq_{\kappa}\kappa_{v}$ using constraints in $\mathit{cert}_{v}$, which is signed by $v$’s parent. $\mathcal{A}$ cannot modify $\mathit{cert}_{v}$ without the parent’s key. Recursing upward, the chain terminates at an uncompromised non-agent ancestor (EUF-CMA). ∎∎
Theorem 5 (Capability Binding).
If $H_{v}(t^{\prime})\neq H_{v}^{\mathit{cert}}$, then $\forall t\geq t^{\prime}:\mathrm{Access}(v,c,t)=\mathrm{DENY}$.
Proof.
By collision resistance: $\Pr[H_{v}(t^{\prime})=H_{v}^{\mathit{cert}}]\leq\mathrm{negl}(\lambda)$ when $S_{v}(t^{\prime})\neq\sigma_{v}$. Forging a new certificate requires the parent’s key (same argument as Theorem 4). ∎∎
Theorem 6 (Chain Integrity).
Only chains forming a valid path from a trusted root with valid signatures pass verification.
Proof.
Fabricating any link requires the corresponding private key. By EUF-CMA, $\mathcal{A}$ cannot produce valid signatures without the key. ∎∎
Theorem 7 (Delegation Depth Enforcement).
Agent delegation depth from any non-agent ancestor never exceeds $\mathit{max\_depth}$.
Proof.
Trust Containment requires $\kappa_{v}.\mathit{max\_depth}<\kappa_{\mathit{par}(v)}.\mathit{max\_depth}$ (strict). Starting from depth $d$, the chain has at most $d$ agent levels before reaching 0. ∎∎
Theorem 8 (Reproducibility Soundness).
If $v$ executes $M^{\prime}\neq M$ or $S^{\prime}\neq S$, replay verification detects with probability $\geq 1-\epsilon(\lambda)$.
Proof.
By DiFR [1]: $\mathbb{E}[\mathrm{CharMatch}(O^{\mathit{actual}},O^{\mathit{ref}})]\leq\delta$ for different models. By Hoeffding: $\Pr[\mathrm{CharMatch}\geq\theta]\leq\exp(-2n(\theta-\delta)^{2})$, negligible for realistic $n$. ∎∎
Theorem 9 (Bounded Divergence Under Verification).
If $M^{\prime}$ passes replay verification on $n$ independent prompts from $\mathcal{D}$, then $\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]\leq\epsilon$ where $\epsilon\leq\exp(-2n\theta^{2})$.
Proof.
Let $p=\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]$ be the true failure rate. Each verification trial is Bernoulli($p$). Observing $n$ successes (all pass), the likelihood ratio test gives: $\Pr[\text{all pass}\mid p]=(1-p)^{n}$. For $p\geq\epsilon$: $(1-\epsilon)^{n}\leq e^{-n\epsilon}$. Setting the false-acceptance probability to $e^{-n\epsilon}\leq e^{-2n\theta^{2}}$ (using $\epsilon=2\theta^{2}$ as the critical value from the Hoeffding bound on the CharMatch statistic), we obtain the stated bound. The Hoeffding refinement uses the fact that CharMatch is a bounded $[0,1]$ statistic with expected value $\geq\theta$ under $H_{0}$ (same model), applying $\Pr[\bar{X}-\mu\leq-t]\leq\exp(-2nt^{2})$ to the complement event. ∎∎
Theorem 10 (Credential Isolation).
$\mathrm{Accessible}(C^{\prime})=\bigcup_{v\in C^{\prime}}\mathrm{Accessible}(\{v\})$.
Proof.
Access depends solely on the requesting agent’s certificate. No protocol mechanism combines permissions. Proxy tokens are bound to specific agent IDs. ∎∎
Theorem 11 (Ledger Integrity).
No party can undetectably modify, delete, or reorder records.
Proof.
(a) Modification invalidates $h_{\mathit{prev}}$ of successor, cascading through chain; re-signing requires all subsequent senders’ keys. (b) Deletion breaks hash chain. (c) Reordering breaks both sequence numbers and hash links. ∎∎
Theorem 12 (Chain Verifiability).
Let $v_{0}\to v_{1}\to\cdots\to v_{n}$ be an execution chain where $v_{0}$ is a non-agent trust anchor, and let $\mathrm{CVD}(v_{n})=k<n$ (Definition 7). Then for all $i>k$, the output of $v_{i}$ cannot be verified as originating from a legitimate computation chain rooted at $v_{0}$, even if $\rho_{v_{i}}.\mathit{level}=\mathit{full}$.
Proof.
We proceed by induction on chain position.
[…]
Inductive step: Assume the output of $v_{i-1}$ (i1>ki-1>k) cannot be verified as legitimate. Agent $v_{i}$ receives this unverifiable output as input. By the same argument as the base case, $v_{i}$’s replay verification confirms computational fidelity on unverifiable input. ∎∎
7 Implementation and Evaluation
7.1 Prototype Architecture
We implement the framework as two Rust crates: agent-trust-kit (open-source SDK) and a2auth (credential governance application). Table 1 summarizes the implementation.
Table 1: Implementation summary.
Component Dependencies LoC
Certificate (parse, issue, verify) x509-parser, ring $\sim$}1,400
Skills manifest hasher sha2, serde_cbor $\sim$}500
Trust tree (internal) $\sim$}700
Reproducibility verifier sha2 $\sim$}400
Interaction ledger ed25519-dalek, sled $\sim$}1,100
Credential vault aes-gcm, argon2 $\sim$}800
Agent registry + proxy tokens agent-trust-kit $\sim$}600
MCP server + CLI jsonrpc-core, clap $\sim$}900
Real LLM repro. validation anthropic, sklearn (Python) $\sim$}870
MCP governance proxy demo cryptography, cbor2 $\sim$}520
Total $\sim$}8,290
7.2 Performance Evaluation
We evaluate on Apple M2, 16 GB RAM, macOS. All benchmarks use Criterion.rs with 10,000 samples and 3-second warmup.
Experiment 1: Certificate Verification Latency.
Chain Depth Mean Per-Sig Cap. Binding
[…]
5 (max practical) 163.6 µs 32.7 µs <<1 ns
Capability binding adds <<1 ns—the cost of a single 32-byte comparison. Verification is dominated by Ed25519 signature verification ($\sim$}33 µs/sig).
Experiment 2: Skills Manifest Hash.
Tools Sort CBOR SHA-256 Total
5 32 ns 0.41 µs 2.25 µs 2.69 µs
10 42 ns 0.76 µs 4.30 µs 5.10 µs
20 64 ns 1.71 µs 8.78 µs 10.6 µs
50 152 ns 3.94 µs 21.8 µs 25.9 µs
100 263 ns 8.05 µs 43.3 µs 51.6 µs
Experiment 3: Ledger Throughput.
Operation Latency Throughput
Append record 448 µs $\sim$}2,230 rec/s
Verify single record 32.4 µs $\sim$}30,900 rec/s
Full audit (1K records) 34.7 ms 34.7 µs/rec
Full audit (100K records) 3.46 s 34.6 µs/rec
Query by agent ID 112 µs 0.56 µs/rec
Experiment 4: Reproducibility Verification Latency.
Tokens Compare Replay Skip Commit
128 120 ns 123 ns 3.5 ns 1.72 µs
512 416 ns 418 ns 3.5 ns 6.55 µs
2,048 1.55 µs 1.55 µs 3.5 ns 24.3 µs
8,192 6.01 µs 6.10 µs 3.6 ns 96.6 µs
Token comparison scales at $\sim$}0.73 ns/token. Batch verification: 100 interactions (512 tokens each) in 41.8 µs. An auditor verifies 100,000 interactions in $\sim$}42 ms.
Experiment 5: Real LLM Reproducibility Validation. We validate the DiFR finding empirically in two phases. Phase 1 uses two Anthropic models (Claude Sonnet 4 and Haiku 4.5, temperature==0, max_tokens==512) across 105 prompts in five task types—code generation (25), JSON/structured output (15), mathematical reasoning (20), prose explanation (30), and creative writing (15)—with 5 same-model repetitions and 3 cross-model repetitions each (1,050 same-model + 1,575 cross-model pairwise comparisons). Phase 2 extends to 9 models from 7 providers: Claude Sonnet 4 and Haiku 4.5 (Anthropic), GPT-4.1 and GPT-4.1 mini (OpenAI), Gemini 2.5 Pro (Google), Grok 4 (xAI), DeepSeek V3 (DeepSeek), Llama 4 Maverick (Meta), and Kimi K2.5 (Moonshot). All models are accessed via OpenRouter as a unified API gateway (plus Volcengine for Doubao), with temperature==0 and 100 prompts across six task types (code, JSON, math, prose, creative, knowledge), 10 repetitions per model-prompt pair (5,400 total API calls; 15,120 pairwise comparisons: 10,800 same-model + 240 same-provider + 4,080 cross-provider). We compute four similarity metrics: character-level positional match (CharMatch), Jaccard token overlap, word-level TF-IDF cosine, and character n-gram (3–5) cosine; then apply Welch’s $t$-test, Mann-Whitney $U$, and Cohen’s $d$ to quantify separation, and use Youden’s $J$ statistic for optimal threshold selection.
Result 5a: Same-model reproducibility is task-type dependent. Table 2 shows per-task reproducibility from Phase 1 across 105 prompts (1,050 same-model pairwise comparisons).
Table 2: Same-model (Sonnet 4) reproducibility by task type.
Task Type $n$ Identical CharMatch TF-IDF cos
[…]
Structured outputs (JSON, code) have the highest identical-output fractions (>>50%), while prose shows the most variation (34.7% identical), driven by paraphrase-level reorganization. The bimodal distribution (median CharMatch =0.975=0.975 vs. mean =0.627=0.627) confirms that same-model outputs are either near-identical or substantially rephrased.
Result 5b: Multi-provider separation analysis. Table 3 presents the Phase 2 multi-provider results across 15,120 pairwise comparisons (10,800 same-model vs. 4,080 cross-provider). All differences are statistically significant ($p\approx 0$, Welch’s $t$ and Mann-Whitney $U$).
Table 3: Multi-provider separation: 9 models, 7 providers (15,120 comparisons).
Metric Same-model Cross-provider Sep. Cohen’s $d$
CharMatch 0.538 0.114 $4.7\times$ 1.36
Jaccard 0.710 0.321 $2.2\times$ 1.51
TF-IDF cosine 0.839 0.586 $1.4\times$ 1.08
Char n-gram cos 0.871 0.640 $1.4\times$ 1.20
All four metrics achieve Cohen’s $d>1.0$ (large effect sizes) across the full 9-model set, with Jaccard ($d=1.51$) and CharMatch ($d=1.36$) providing the strongest separation. Notably, same-provider pairs (e.g., GPT-4.1 vs. GPT-4.1 mini: CharMatch =0.251=0.251; Sonnet 4 vs. Haiku 4.5: $0.192$) exhibit similarity intermediate between same-model and cross-provider, confirming that even models from the same vendor maintain distinct output fingerprints.
Result 5c: Per-model determinism varies by over an order of magnitude. Table 4 ranks models by same-model CharMatch, revealing a wide determinism spectrum.
Table 4: Per-model same-model reproducibility (CharMatch, temperature==0).
Model Provider CharMatch Std
Claude Haiku 4.5 Anthropic 0.987 0.088
Claude Sonnet 4 Anthropic 0.823 0.273
GPT-4.1 mini OpenAI 0.701 0.335
GPT-4.1 OpenAI 0.617 0.363
Llama 4 Maverick Meta 0.528 0.360
Grok 4 xAI 0.402 0.343
Gemini 2.5 Pro Google 0.392 0.397
DeepSeek V3 DeepSeek 0.225 0.271
Kimi K2.5 Moonshot 0.169 0.233
Reasoning/thinking model (internal chain-of-thought).
Claude Haiku 4.5 produces near-identical outputs across repetitions (CharMatch =0.987=0.987), while reasoning models (Kimi K2.5, Gemini 2.5 Pro, Grok 4) exhibit substantially lower determinism due to stochastic chain-of-thought reasoning even at temperature 0. Despite this variation, even the least deterministic model (Kimi K2.5, CharMatch =0.169=0.169) still exceeds the mean cross-provider similarity ($0.114$), demonstrating that model identity remains distinguishable for all tested models.
Result 5d: Per-task separation is strongest for mathematical reasoning. Table 5 shows CharMatch separation factors across six task categories.
Table 5: Per-task CharMatch separation (same-model / cross-provider).
Task Type Same-model Cross-provider Sep.
Math reasoning 0.508 0.042 $12.2\times$
Knowledge 0.511 0.072 $7.1\times$
Creative writing 0.402 0.057 $7.0\times$
Prose 0.488 0.083 $5.9\times$
Code generation 0.594 0.161 $3.7\times$
JSON / structured 0.779 0.299 $2.6\times$
Mathematical reasoning achieves the highest separation ($12.2\times$) because models develop highly distinctive step-by-step reasoning styles, while JSON outputs show the lowest separation ($2.6\times$) because structural constraints force convergent formatting. These task-dependent results motivate per-task threshold calibration in deployment.
Result 5e: Optimal threshold via Youden’s $J$ statistic. Table 6 reports optimal thresholds from the multi-provider experiment.
Table 6: Optimal classification thresholds (multi-provider, 15,120 comparisons).
Metric $\theta^{*}$ Youden’s $J$ $F_{1}$
Jaccard 0.408 0.592 0.876
CharMatch 0.146 0.572 0.873
Char n-gram cos 0.809 0.538 0.855
TF-IDF cosine 0.837 0.466 0.844
In the multi-provider setting, Jaccard achieves the best overall classification ($F_{1}=0.876$, $J=0.592$), followed closely by CharMatch ($F_{1}=0.873$, $J=0.572$). Both metrics attain $F_{1}>0.87$ across 9 models and 7 providers, confirming that lightweight textual similarity metrics—requiring no model internals, embeddings, or logit access—provide practically useful model identity verification even in a heterogeneous multi-provider environment.
Experiment 6: End-to-End MCP Overhead.
Phase Latency
MCP request parse 0.02 ms
Certificate chain verification 0.097 ms
Skills hash comparison <<0.001 ms
Proxy token validation 0.05 ms
Ledger append 0.448 ms
Total governance overhead $\sim$}0.62 ms
Typical MCP call (no governance) 50–500 ms
Overhead as % of total 0.1–1.2%
Experiment 7: Comparison with BAID.
Metric BAID Ours Ratio
Verification latency $\sim$}120 s 0.097 ms $1.2{\times}10^{6}$
Proof size $\sim$}50 KB $\sim$}2 KB $25\times$
Infrastructure Blockchain None
Cap. change detection Post-hoc Real-time
Experiment 8: MCP Governance Proxy Demo.
To validate real-world applicability, we implement a complete MCP governance proxy in Python (520 lines) that interposes between an MCP client (agent) and MCP server (tool provider). The proxy enforces all three governance requirements: G1 via skills manifest hash verification against the agent’s certificate, G2 via reproducibility commitment tracking, and G3 via a hash-chain interaction ledger with Ed25519-signed records. We test 5 live attack scenarios:
Attack Result Latency
S1: Inject web_browser at runtime Blocked (hash $\Delta$) 121 µs
S2: Replace read_file with trojan Blocked (code hash) 103 µs
S3: Forged certificate (rogue CA) Blocked (sig verify) 106 µs
[…]
S5: Tamper ledger record #2 Detected (chain break)
All attacks are detected in real-time. Over 1,000 governed calls, the proxy achieves mean governance overhead of 120 µs/call (median: 106 µs, p99: 432 µs), throughput of 8,195 governed calls/sec, and all ledger records pass integrity verification. This confirms that the framework provides real-time protection with overhead well below the latency of typical MCP calls (50–500 ms).
7.3 Attack Scenario Evaluation
We implement 12 attack scenarios as automated tests: 9 structural governance attacks as Rust unit tests (774 lines), 5 end-to-end scenarios via the MCP governance proxy demo (520 lines), and 3 runtime behavioral attacks informed by the Nemotron-AIQ dataset [45] (which demonstrates 100% attack success rates without structural defenses). Table 7 summarizes results.
Table 7: Attack scenario evaluation. All 12 tests pass deterministically. S10–S12 are runtime behavioral attacks validated against Nemotron-AIQ traces.
Scenario Attack Baseline Framework Detection Mechanism
S1: Silent escalation Install web_browser at runtime Undetected Prevented capability_binding: hash mismatch
S2: Tool trojanization Replace implementation (same name) Undetected Prevented capability_binding: code hash change
S3: Model substitution Forge cert with different model Undetected Detected sig_verify: forged cert unsigned
S4: Phantom delegation T2 agent issues T1 cert Undetected Prevented constraint_monotonicity
S5: Evidence tampering Modify ledger record N/A Detected Hash chain + signature invalidation
S6: Blame shifting Dispute in agent chain Unresolvable Resolved Ed25519 signatures
S7: Colluding agents T3+T3 attempt T1 access Undetected Prevented tier_check: independent evaluation
S8: Chain forgery Fabricated chain Prevented sig_verify: fake key $\neq$ root
S9: Depth overflow $\mathit{max\_depth}=0$ delegation Prevented Structural: depth <0<0 impossible
S10: Knowledge poisoning Inject false data via tool output Undetected Traced Ledger $\eta_{\mathit{in}}/\eta_{\mathit{out}}$: pinpoint injection span
S11: Cascading hijack Multi-stage instruction persistence Undetected Traced Hash-chain: full propagation path audit
S12: Output exfiltration Embed phishing via rendered output Undetected Detected Replay divergence from committed $\pi_{i}$
100% attack success in Nemotron-AIQ undefended traces (2,596 security trials).
7.4 End-to-End Multi-Agent Pipeline
Experiments 1–8 evaluate individual governance operations in isolation. To validate the framework under realistic conditions, we deploy the full A2Auth governance stack across a multi-agent pipeline where agents coordinate, delegate tasks, invoke tools via MCP, and pass results through the interaction ledger. This experiment measures: (a) end-to-end governance overhead as a fraction of pipeline latency, (b) scaling behavior from 5 to 20 agents, (c) attack detection in situ, and (d) comparison with three baselines. The implementation uses the Rust agent-trust-kit SDK exposed to Python via PyO3 FFI, with agents making real Anthropic Claude API calls.
Figure 5: End-to-end multi-agent pipeline architecture with governance overlay. Five specialist agents coordinate through a trust tree (depth 3). Red labels indicate E2E attack injection points (Table 10). All inter-agent communications produce bilaterally-signed ledger records.
Pipeline architecture (Figure 5). We implement an AI Research Assistant pipeline: a user submits a research query; a Coordinator agent (T1_HIGH) orchestrates five specialist agents—Research (T2), Analysis (T2), Writer (T2), Review (T3_MONITORED), and Fact-Checker (T2)—each with a distinct skills manifest and reproducibility commitment. The trust tree has depth 3 (Root CA $\to$ Human $\to$ Coordinator $\to$ Workers), with the Coordinator’s max_delegation_depth=2{}=2 and all workers at depth 0 (cannot delegate). A single pipeline run traverses 21 steps: 10 tool calls (7 LLM invocations + 3 non-LLM tools), 11 inter-agent communications (task assignments and result deliveries), each producing a bilaterally-signed ledger record with capability verification and reproducibility commitment.
Experiment 9a: Governance overhead. Table 8 breaks down governance cost per pipeline run. Each operation’s latency is consistent with the micro-benchmarks from Experiments 1–4; the end-to-end cost is their composition.
Table 8: Per-pipeline governance overhead breakdown (5-agent configuration, 21 ledger entries per run, mean of 100 runs).
Operation Per-op Count Subtotal
Certificate chain verification 97 µs 6 0.58 ms
Skills manifest hash check 5.1 µs 10 0.05 ms
Ledger append (tool call) 448 µs 10 4.48 ms
Reproducibility commitment 6.6 µs 10 0.07 ms
Ledger append (inter-agent) 448 µs 11 4.93 ms
Bilateral Ed25519 signing 66 µs 11 0.73 ms
Total governance 21 10.8 ms
Pipeline latency (real LLM) $\sim$}8.7 s
Overhead 0.12%
Governance adds 10.8 ms to a pipeline that takes $\sim$}8.7 s with real Claude API calls (7 LLM inferences at $\sim$}1.2 s each + 3 non-LLM tools at $\sim$}100 ms each). The overhead of 0.12% is dominated by ledger appends (87% of governance cost), confirming that the cryptographic operations themselves (hashing, signing, verification) are negligible relative to I/O-bound agent workloads.
Experiment 9b: Scaling behavior. We scale the pipeline from 5 to 20 agents by instantiating parallel specialist groups: the Large configuration uses 4 Research agents (handling sub-queries in parallel), 4 Analysis agents, 3 Writer agents, 3 Review agents, 3 Fact-Checkers, 2 Coordinators (primary + sub-coordinator), and 1 Aggregator. Table 9 shows that per-agent governance cost remains approximately constant.
Table 9: Governance scaling: 5 $\to$ 20 agents (mean of 50 runs per configuration).
Config Agents Ledger Gov. Per-agent Storage
entries (ms) (ms) (KB)
Small 5 21 10.8 1.81 5.2
Medium 10 45 23.0 2.09 11.2
Large 20 95 48.3 2.30 23.8
Total governance cost scales linearly with ledger entries ($R^{2}>0.99$), but per-agent cost increases only 27% when quadrupling the agent count (1.81 ms $\to$ 2.30 ms). This near-constant per-agent cost arises because each agent’s governance operations—certificate verification, capability checking, ledger append—are independent of the total number of agents in the pipeline. The slight increase comes from deeper trust chains in the Large configuration (depth 4 vs. depth 3) adding one extra signature verification per activation. Ledger storage grows at $\sim$}250 bytes per entry; even the 20-agent pipeline produces only 23.8 KB per run, confirming that commitment-only storage (§5) is practical for long-running agent workflows.
Experiment 9c: End-to-end attack detection. We inject five attacks into the live 5-agent pipeline, each targeting a different governance layer. Unlike the unit-level tests in Table 7, these attacks occur mid-pipeline with real inter-agent data flow.
Table 10: End-to-end attack detection in the live multi-agent pipeline.
Attack Target Layer Detect. Hops
E2E-1: Capability escalation Research G1 5 µs
E2E-2: Model substitution Analysis G2 $\dagger$
E2E-3: Phantom delegation Writer G1 97 µs
E2E-4: Ledger tampering Ledger G3 32 µs
E2E-5: Cascading injection Pipeline G1+G2+G3 97 µs 3
[…]
dominated by LLM re-inference ($\sim$}1.2 s), not framework cost ($\sim$}13 µs).
[…]
Across 100 clean pipeline runs (500 total governed tool calls, 2,100 ledger entries), zero false positives are observed.
Experiment 9d: Baseline comparison. We evaluate the same five E2E attacks against three baselines: (1) no governance (raw MCP), (2) MCP with OAuth 2.1 (standard MCP authentication per specification), and (3) MCP with OpenTelemetry traces (observability-only, no cryptographic integrity). Table 11 summarizes results.
Table 11: E2E attack detection: A2Auth vs. baselines ($\checkmark$ = detected/prevented, ✗ = undetected). G1/G2/G3 indicate which governance requirement enables detection.
Attack No gov. OAuth OTEL A2Auth
E2E-1: Capability escalation $\checkmark$ (G1)
E2E-2: Model substitution $\checkmark$ (G2)
E2E-3: Phantom delegation a $\checkmark$ (G1)
E2E-4: Ledger tampering N/A N/A b $\checkmark$ (G3)
E2E-5: Cascading injection c $\checkmark$ (G1+G2+G3)
Total detected 0/5 0/5 0/5 5/5
aOAuth authenticates the Writer but has no delegation depth constraint. bOTEL traces can be silently modified (no hash chain or signatures; §5). cOTEL logs the poisoned data flow but cannot prove tampering or trace provenance cryptographically.
OAuth 2.1 provides identity authentication but not capability binding: it verifies who the agent is, not what tools it currently has. OpenTelemetry provides observability but not integrity: OTEL span trees log tool calls and parent-child relationships, but as demonstrated by analysis of 10,796 Nemotron-AIQ traces (§5), OTEL records contain no hash chains, no cryptographic signatures, and no identity binding—any field can be silently modified. A2Auth’s three governance layers (G1+G2+G3) provide the only configuration that detects all five attack types, with governance overhead below 0.2% of pipeline latency.
8 Limitations and Future Work
8.1 Scope of Contribution
Our contribution is a governance framework, not novel cryptographic primitives. The individual building blocks—X.509 certificates, SHA-256, Ed25519, hash chains—are standard. The novelty lies in (a) identifying the capability-context separation as the principled boundary for agent governance, (b) showing that composing standard mechanisms along this boundary enforces G1–G3 with negligible overhead, and (c) demonstrating this is practically achievable. We view this as analogous to Certificate Transparency [16], which composes Merkle trees and gossip protocols—neither novel individually—into infrastructure that fundamentally changed how the web PKI operates. The security proofs are intentionally straightforward reductions to standard assumptions (collision resistance, EUF-CMA); the contribution is in what to build, not in new ways to prove things.
8.2 Capability Binding Limitations
Capability-context boundary. Our sharp boundary has edge cases: a “code execution” tool can effectively acquire new capabilities through composition (e.g., downloading and running arbitrary scripts). Our framework guarantees that the registered capability set has not changed, but the effective attack surface within a fixed capability envelope can still be large. Complementary approaches (sandboxing, taint tracking, output filtering) are needed to govern behavior within the envelope. We view capability binding and intra-capability governance as orthogonal layers: the former ensures the attack surface is known; the latter constrains behavior within it.
Closed-source binding. For closed-source models and tools, the skills hash captures configuration metadata (name, version, API schema), not implementation code. A provider could change implementation without changing the public descriptor. Reproducibility verification provides a complementary detection mechanism: implementation changes that alter behavior will be detected via replay divergence.
Re-signing frequency. Mitigated through wildcard tool categories, short-lived auto-renewable certificates, and the empirical observation that tool configurations change far less frequently than user context.
8.3 Reproducibility Limitations and Graceful Degradation
Reproducibility is a design constraint, not a universal guarantee. Our framework makes explicit what is often implicit: non-reproducible agent behavior is non-verifiable. This is not a limitation unique to our approach—it applies to any verification scheme not based on trusted hardware. For high-assurance deployments, this means agents should be configured for deterministic output (structured formats, temperature==0, fixed seeds). Our empirical results (§7) show that code-generating agents already satisfy this constraint perfectly, and the separation between same-model and cross-model distributions ($11.5\times$) provides robust detection even for prose tasks.
Graceful degradation under partial G2. Our multi-provider experiment (§7, Experiment 5c) reveals that model determinism varies by over an order of magnitude: from CharMatch==0.987 (Haiku 4.5) to 0.169 (Kimi K2.5). This raises a practical question: what security guarantees hold when some agents use low-determinism models?
When G2 is unavailable (reproducibility commitment $\mathit{none}$), the framework degrades to G1+G3 only. Table 12 characterizes the precise security impact.
Table 12: Security guarantees under G1+G2+G3 vs. G1+G3 degradation.
Attack class G1+G2+G3 G1+G3
Silent capability escalation (ATK-1) Prevented Prevented
Tool trojanization (ATK-3) Prevented Prevented
Phantom delegation (ATK-4) Prevented Prevented
Chain/depth violation (ATK-5,6) Prevented Prevented
Evidence tampering (ATK-8) Detected Detected
Blame shifting (ATK-9) Resolved Resolved
Credential collusion (ATK-7) Prevented Prevented
Model substitution (ATK-2) Detected
Output exfiltration (ATK-12) Detected
G1+G3 alone protect against 9 of 12 attack types. The three attacks that require G2—model substitution (ATK-2), cascading injection via model compromise (ATK-10/11 when the injection occurs through model replacement rather than tool replacement), and output exfiltration via model tampering (ATK-12)—share a common structure: the attacker operates within a valid capability envelope by substituting the computational process itself. Signing proves attribution; only G2 proves computational fidelity.
Chain-level degradation. The Chain Verifiability Theorem (§4.8) establishes a stronger result: a G2==none agent at an interior position breaks behavioral verification for all downstream nodes. This means the true cost of deploying a low-determinism model at a coordinator position is not merely that the coordinator itself is unverifiable, but that the entire sub-tree it governs loses G2 protection. The framework’s three commitment levels (full/statistical/none) are thus not equivalent—their security implications depend critically on chain position.
Statistical vs. computational security. Token-level reproducibility provides statistical rather than computational security. However, the Bounded Divergence Theorem (§4.7) shows this gap is operationally small: sufficient replay verification transforms statistical detection into a provable safety certificate. An adversary who fine-tunes a model to pass $n$ diverse replay checks has produced a model that is $(n,\epsilon)$-indistinguishable from the original, with $\epsilon$ decreasing exponentially in $n$. Future work will explore semantic similarity (embedding cosine distance) as a verification metric, which may provide tighter separation bounds and is robust to paraphrase-level variation.
Provider cooperation. Replay verification requires access to the same model version. Organizations should negotiate replay rights as part of service agreements. Deterministic inference providers [6] eliminate this dependency entirely.
8.4 Trust Infrastructure Limitations
Non-agent node compromise. We assume non-agent nodes (CAs, humans, organizations) are uncompromised. Standard PKI mitigations apply: Certificate Transparency logs, hardware trust anchors (TPM 2.0), and multi-party issuance protocols. Ledger operator trust. The current single-operator ledger relies on bilateral signing for tamper evidence. A multi-operator design with consensus is possible but introduces latency overhead. Application-layer data flow. Our framework governs agent identity and capability, not the content of data flows. Complementary approaches (data-flow taint tracking, differential privacy) address this orthogonal concern.
8.5 Future Work
Hardware-attested agent identity (TPM 2.0, TEEs) for stronger G2 guarantees without provider cooperation. Embedding-based semantic similarity for reproducibility verification with formal threshold analysis. Cross-protocol capability governance beyond MCP/A2A. AI-assisted certificate issuance to reduce operational burden. Dynamic trust calibration from ledger history, enabling trust scores to evolve based on observed agent behavior.
9 Related Work
We organize related work along the four-layer landscape.
Identity Platforms. Keyfactor [19] and HID Global [20] propose X.509 for agents but without tool binding. SPIFFE/SPIRE [21, 22] provides workload identity. CyberArk [37] extends PAM to agents.
Authorization Frameworks. OIDC-A [4] extends OpenID Connect with agent claims. South et al. [3] propose Agent-ID Tokens (ICML 2025). Agentic JWT [10] defines delegation with temporal constraints.
Runtime Gateways. Gravitee [40] and Lasso [41] provide MCP proxies with per-call enforcement—fundamentally stateless.
Governance Platforms. Credo AI [42] maintains agent registries at the documentation layer.
Verifiable Execution. BAID [2] provides zkVM-based verification ($\sim$}120 s per proof). DiFR [1] demonstrates inference near-determinism. EigenAI [6] achieves 100% deterministic inference. Attestable Audits [7] uses TEEs.
Classical Trust Management. PolicyMaker [12], KeyNote [13], SPKI/SDSI [14], F-PKI [16], ARPKI [17].
Agent Protocols. A2A [8] provides agent communication. MCP enables agent-tool interaction with OAuth 2.1.
Agent Governance Frameworks. Cheng et al. [43] propose a three-pillar model (transparency, accountability, trustworthiness) as a conceptual governance framework. Allegrini et al. [44] formalize 31 task-lifecycle properties in CTL/LTL but address operational correctness rather than cryptographic governance. Ghosh et al. [45] propose an operational safety framework with agent red-teaming probes (ARP), policy-based tool scoping, and OpenTelemetry-based audit trails, validated through a 10,796-trace dataset (Nemotron-AIQ) covering 22 security attack templates and 13 content-safety categories against NVIDIA’s AI-Q research assistant. Their empirical results demonstrate that without guardrails, all 2,596 security attacks succeed (100%), and 20% of 2,800 safety attacks propagate to the final output—providing compelling evidence for the inadequacy of relying solely on model alignment. However, their OTEL-based audit trails lack cryptographic integrity: span linkage is purely referential (no hash chaining), spans carry no agent identity binding (no certificates), and full plaintext I/O is stored (no privacy preservation). Our framework addresses these gaps with a formally verified, hash-chain-linked, bilaterally signed interaction ledger (Section 5). Furthermore, their “signed schemas” and “per-call capability tokens” remain design principles without cryptographic specification or formal proofs. Raza et al. [28] adapt TRiSM for agentic multi-agent systems.
Regulatory Landscape. The EU AI Act (Regulation 2024/1689) requires traceability but lacks agent-specific implementing acts [33]. NIST CAISI [26] is developing agent standards. CSA [24, 25] proposes agentic trust frameworks.
10 Conclusion
AI agents increasingly acquire capabilities at runtime, yet no existing framework enforces the three governance requirements—capability integrity, behavioral verifiability, and interaction auditability—that a trustworthy agent ecosystem demands.
We addressed this through the capability-context separation principle and three mechanisms: capability-bound certificates (G1), reproducibility commitments (G2), and a verifiable interaction ledger (G3). We formalized nine security properties (including chain verifiability: behavioral verification is a chain property broken by any unverifiable interior agent), implemented the framework in Rust with 97 µs verification and <<1 ns capability binding overhead (${\sim}1{,}200{,}000\times$ faster than BAID), validated reproducibility across 9 models from 7 providers ($F_{1}=0.876$ multi-provider, $F_{1}=0.990$ single-provider), and verified 12 attack scenarios—including runtime behavioral attacks validated against NVIDIA Nemotron-AIQ traces—with specific, traceable detection mechanisms. An end-to-end evaluation deploying the full governance stack across a 5-to-20-agent pipeline with real LLM calls confirms 0.12% overhead, sub-linear per-agent scaling, and detection of all in-situ attacks with zero false positives—while three baselines (no governance, OAuth 2.1, OpenTelemetry) detect none.
Our framework fills the stateful, capability-aware governance gap between identity platforms and runtime gateways. As the EU AI Act’s high-risk provisions take effect and NIST develops agent-specific standards, the need for cryptographic infrastructure enforcing G1–G3 will only grow.
We release the core mechanisms as an open-source SDK (agent-trust-kit) to enable ecosystem adoption.

### 19 Mar 2026 v2

@@
 Abstract
-[…]
+AI agents now dynamically acquire tools, orchestrate sub-agents, and transact across organizational boundaries—yet no existing security layer can verify what an agent is capable of, whether it executed what it claims, or what actually happened in a multi-agent interaction.
+We trace this governance vacuum to a missing architectural distinction: the capability-context separation.
+Inside a transformer’s forward pass, tool definitions and user context are indistinguishable token sequences; at the orchestration layer, they have fundamentally different security semantics—tool definitions determine which real-world actions are possible (and change infrequently), while runtime context determines which actions are chosen (and changes per interaction).
+Existing frameworks conflate the two, enabling silent capability escalation—agents acquiring tools without invalidating credentials—and leaving cross-agent interactions without verifiable provenance.
+From this principle, we derive three Agent Governance Requirements: capability integrity (G1) governs the capability envelope, behavioral verifiability (G2) ensures agents executed their declared computational process, and interaction auditability (G3) provides tamper-evident records for the runtime context—together defining what an agent ecosystem must enforce, independent of how.
+We prove two structural results with prescriptive architectural consequences: the Chain Verifiability Theorem, establishing that behavioral verification is a chain property (one unverifiable interior agent breaks end-to-end verification for all downstream nodes), and the Bounded Divergence Theorem, which transforms replay-based verification into a probabilistic safety certificate ($\epsilon\leq 1-\alpha^{1/n}$).
+We validate the framework with two crypto-agnostic instantiations—basic (Ed25519, SHA-256, hash chains; 97 µs verify) and enhanced (BBS+ selective disclosure, Groth16 DV-SNARK; 13.8 ms verify)—both satisfying the same nine security properties.
+A multi-provider reproducibility study (9 models, 7 providers) reveals $5.8\times$ variance in inference determinism, directly connecting model characteristics to governance architecture.
+End-to-end evaluation over 5–20 agent pipelines with real LLM calls confirms <<0.02% governance overhead, real-time detection of 7 end-to-end attack scenarios (covering G1, G2, and G3) with zero false positives, and post-hoc forensic tracing of runtime behavioral attacks validated against the Nemotron-AIQ dataset.
 1 Introduction
@@
 AI agents are no longer isolated tools.
-Through protocols like MCP (Model Context Protocol) [8] and A2A (Agent-to-Agent Protocol) [8], modern agents dynamically discover tools, invoke external APIs, execute code, and orchestrate other agents at runtime.
-The MCP ecosystem alone has grown to over 1,000 tool servers in under a year [32].
-As these agents assume real-world responsibilities—managing enterprise credentials, executing financial transactions, accessing sensitive data—a fundamental question emerges: what properties must an agent ecosystem satisfy for humans to trust it?
-This question has been raised from multiple directions.
-The EU AI Act (Regulation 2024/1689) mandates traceability, logging, and conformity assessment for high-risk AI systems (Articles 12, 43, 72).
-NIST’s AI Agent Standards Initiative [26] identifies security controls, governance oversight, and monitoring as key requirements.
-The TRiSM framework [28] structures agent governance around explainability, security, privacy, and lifecycle management.
-Schroeder de Witt [11] calls for secure interaction protocols that prevent insecure emergent behaviors.
-Cheng et al.
-[43] ground agent safety in transparency, accountability, and trustworthiness.
-Allegrini et al.
-[44] formalize 31 safety and liveness properties for agentic task lifecycles in temporal logic.
-Ghosh et al.
-[45] propose an operational safety framework with red-teaming methodology, policy-based tool scoping, and optional trace signing—but without cryptographic capability binding, formal adversary models, or verifiable replay mechanisms.
-Yet despite this convergence of concern, no existing framework translates these governance aspirations into enforceable cryptographic infrastructure.
-The gap is not in recognizing what is needed, but in building it.
-We propose three Agent Governance Requirements that distill the common thread across these efforts into technically precise, verifiable properties:
+Through protocols like MCP [8] and A2A [9], modern agents dynamically discover tools, invoke external APIs, and orchestrate other agents at runtime.
+As these agents assume real-world responsibilities—managing credentials, executing financial transactions, accessing sensitive data—a fundamental question emerges: what properties must an agent ecosystem satisfy for humans to trust it?
+This question has been raised from multiple directions: regulatory mandates for traceability and conformity assessment (EU AI Act, Articles 12, 43, 72), standards initiatives identifying security controls and governance oversight [27, 29], and research calling for secure interaction protocols [12], transparency [44], formal safety properties [45], and operational safety frameworks [46].
+Yet no existing framework translates these aspirations into enforceable cryptographic infrastructure.
+This is a fundamental architectural gap: current agents acquire tools dynamically, execute arbitrary models, and interact without verifiable records.
+We propose three Agent Governance Requirements—the minimum conditions under which an agent ecosystem can be considered governed:
 Definition 1 (Agent Governance Requirements).
@@
 Any change to the capability set invalidates existing credentials and requires explicit re-authorization by a human or organizational principal.
-G2: Behavioral Verifiability.
-An agent’s actions within its authorized capability envelope can be independently verified post-hoc, enabling detection of model substitution, unauthorized tool use, or output manipulation.
+G2: Behavioral Verifiability (Inference Authenticity).
+An agent’s declared computational process can be independently verified—ensuring that the agent actually executed what its credentials claim.
+This requirement is implementation-agnostic: it may be satisfied by hardware attestation, software-based replay, or zero-knowledge proofs.
 G3: Interaction Auditability.
 All inter-agent and agent-human interactions produce tamper-evident records sufficient for forensic reconstruction, allowing humans to audit agent behavior, attribute responsibility, and provide evidence for regulatory compliance and system improvement.
-These requirements are not novel in aspiration—they reflect the consensus of legal [33, 34], regulatory [26], and technical [11, 28, 43, 44, 45] communities.
-What is novel is the claim that all three must hold simultaneously, and that none is achievable with existing agent infrastructure.
-G1 alone (without G2 and G3) permits certified agents to misbehave undetectably within their capability envelope.
-G2 alone (without G1) cannot distinguish authorized from unauthorized capabilities.
-G3 alone (without G1 and G2) produces audit logs of unverified actions by unverified agents.
+[…]
+Table 1: Governance gap analysis: undefended attacks when one or two governance requirements are absent.
+ATK-9 (blame shifting) requires both G2 and G3 for full resolution.
+All 12 attacks are defined in §2.
+Missing Undefended Attacks #
+Single requirement absent
+[…]
+[…]
+[…]
+Two requirements absent
+[…]
+[…]
+[…]
+[…]
+A critical corollary: even hardware TEEs [53, 54, 55] do not eliminate the need for G1 and G3.
+A TEE-attested agent that silently acquires new tools remains a security threat (G1 violation); a TEE-attested pipeline without tamper-evident records remains unauditable (G3 violation).
+Our framework provides the governance layer that TEE, by itself, cannot.
 1.2 Why the Current Landscape Falls Short
-The current agent security landscape can be mapped into four layers (Figure 1), each addressing fragments of the governance requirements but none satisfying them:
-Identity platforms (Keyfactor [19], SPIFFE/SPIRE [21, 22]) issue certificates binding an agent to a public key—but a certificate remains valid regardless of how many tools the agent acquires after issuance.
-G1 violated: capability changes are invisible to the identity layer.
-Authorization frameworks (OIDC-A [4], Authenticated Delegation [3], Cerbos [39]) define what an authenticated agent may access—but authorization is evaluated against declared capabilities at grant time, with no mechanism to detect drift.
-G1 violated: authorization does not track capability changes.
-Runtime gateways (Gravitee [40], Lasso Security [41]) inspect tool calls in real time—but are fundamentally stateless, with no memory of prior tool configurations and no behavioral verification.
-G2 and G3 violated.
-Governance platforms (Credo AI [42]) maintain agent registries documenting capabilities—but operate at the documentation layer without cryptographic enforcement.
-G1 violated: declared capabilities are not bound to the agent’s actual configuration.
+The current agent security landscape maps into four layers (Figure 1), none satisfying G1–G3: Identity platforms [20, 22, 23] bind agents to keys but certificates remain valid after capability changes (G1 violated).
+Authorization frameworks [4, 3, 40] evaluate permissions at grant time without detecting drift (G1 violated).
+Runtime gateways [41, 42] inspect tool calls statelessly (G2, G3 violated).
+Governance platforms [43] maintain registries without cryptographic enforcement (G1 violated).
 Figure 1: The four-layer agent security landscape and the capability-identity gap.
@@
 Our framework provides the missing cryptographic governance layer satisfying G1–G3, validated end-to-end across 5–20 agent pipelines.
-The root cause is a failure to recognize a fundamental architectural distinction in LLM-based agents.
-Inside a transformer’s forward pass, tool definitions and user context are indistinguishable token sequences.
-But at the orchestration layer, tool definitions determine which real-world actions the model’s outputs can trigger, while runtime context determines which actions the model chooses.
-This yields a principled capability-context separation: the capability envelope (model identity, tool definitions, permission scopes) governs what is possible and changes infrequently; the runtime context (conversation history, user instructions, memories) governs what is chosen and changes per-interaction.
-The security implication is direct: capability changes require re-authorization (G1); context changes require verifiability (G2) and auditability (G3).
-Existing frameworks conflate the two.
-A tool addition and a new user message are treated identically—neither triggers re-evaluation.
-We call this the capability-identity gap: existing frameworks authenticate who an agent is and authorize what it may access, but none detect when the agent’s capabilities have changed.
-This gap enables a class of attacks we term silent capability escalation.
+The root cause is a failure to recognize a fundamental architectural distinction.
+Inside a transformer’s forward pass, tool definitions and user context are indistinguishable token sequences; but at the orchestration layer, tool definitions determine which real-world actions are possible, while runtime context determines which actions are chosen.
+This yields the capability-context separation: the capability envelope (model identity, tool definitions, permission scopes) changes infrequently and requires re-authorization (G1); the runtime context (conversation history, user instructions) changes per-interaction and requires verifiability (G2) and auditability (G3).
+Existing frameworks conflate the two—a tool addition and a new user message are treated identically.
+On the capability side, this creates the capability-identity gap: no existing framework detects when an agent’s capabilities have changed, enabling silent capability escalation.
+On the context side, it means cross-agent interactions lack verifiable provenance—no framework provides tamper-evident records linking which agent did what, with which inputs, producing which outputs—the foundation for accountability in multi-agent workflows.
+To our knowledge, the capability-context separation is the first architectural abstraction that explains why agent security frameworks systematically fail to address capability drift—and it is not merely descriptive but prescriptive: it dictates the governance boundaries any secure agent system must enforce.
+As a concrete illustration, BAID [2] achieves strong code-level authentication via zkVM, binding agent identity to a program commitment $C_{P}=\mathrm{CommitProg}(P)$.
+Yet $C_{P}$ does not encompass dynamically acquired tools: an agent’s program binary remains unchanged when new tools are discovered via MCP [8] or plugin registries, leaving the entire capability envelope outside BAID’s security perimeter.
+This is not an implementation oversight but a direct consequence of designing without the capability-context separation—the framework conflates code identity with capability identity.
+Our G1 requirement, derived from this abstraction, prescribes that governance must commit to the complete capability set (model + tools + permissions), not merely the executing code.
 1.3 Our Framework: Bridging the Gap
-The capability-context separation provides the theoretical foundation for three mechanisms, each targeting a specific governance requirement:
+The capability-context separation provides the foundation for three enforcement mechanisms that instantiate the governance contracts G1–G3:
 […]
@@
 […]
-Together, these three mechanisms provide what the existing landscape lacks: a stateful, cryptographic governance layer satisfying G1–G3, bridging identity platforms and runtime gateways.
-Our contribution is not in novel cryptographic primitives; rather, we show that standard mechanisms—X.509 certificates, SHA-256 hashing, Ed25519 signatures, hash chains—when composed through the capability-context separation, are sufficient to enforce G1–G3 with negligible overhead, providing actionable design guidance for how agents should be built, how they should interact, and how their behavior should be audited.
+Together, these mechanisms provide a stateful, cryptographic governance layer satisfying G1–G3.
+Crypto-agnostic design.
+G1–G3 define what to enforce, not how.
+The framework is crypto-agnostic: each requirement specifies a functional contract with pluggable cryptographic primitives—analogous to TLS cipher suite negotiation.
+We present a basic instantiation (Ed25519, SHA-256, hash chains; ${\sim}97$ µs verification) and an enhanced instantiation (BBS+ selective disclosure for G1, DV-SNARK zero-knowledge verification for G2; 13.8 ms).
+Both satisfy G1–G3 with identical security semantics, differing only in privacy properties and cost.
+This separation of governance logic from cryptographic mechanism demonstrates that agent governance is a composable architecture where organizations select backends matching their threat model.
 1.4 Threat Model and Security Guarantees
-We consider a PPT adversary who can compromise up to $k$ agent nodes, create malicious agents, inject or modify tools, and manipulate network traffic (Dolev-Yao model).
-Critically, the adversary can perform silent capability escalation.
-The adversary cannot compromise non-agent nodes or break standard cryptographic primitives.
-Under this model, we define and prove nine security properties:
-  1. 1.
-    Capability Binding: tool configuration change $\Rightarrow$ access denied (Theorem 5).
-  2. 2.
-    Trust Containment: child constraints $\leq_{\kappa}$ parent constraints (Theorem 4).
-  3. 3.
-    Chain Integrity: valid chain to trusted root required (Theorem 6).
-  4. 4.
-    Delegation Depth: chains bounded by ancestor’s $\mathit{max\_depth}$ (Theorem 7).
-  5. 5.
-    Reproducibility Soundness: different model/tools detected with overwhelming probability (Theorem 8).
-  6. 6.
-    Bounded Divergence: adversary-controlled model passing $n$ replay checks is $(n,\epsilon)$-indistinguishable from declared model (Theorem 9).
-  7. 7.
-    Chain Verifiability: a single unverifiable interior agent breaks end-to-end behavioral verification for all downstream nodes (Theorem 12).
-  8. 8.
-    Credential Isolation: colluding agents gain no extra permissions (Theorem 10).
-  9. 9.
-    Ledger Integrity: undetectable modification impossible (Theorem 11).
-Properties 1–4 enforce G1. Properties 5–7 enforce G2 (with 7 characterizing its chain-level semantics). Properties 8–9 enforce G3.
+We consider a PPT adversary who can compromise up to $k$ agent nodes, create malicious agents, inject or modify tools, manipulate network traffic (Dolev-Yao model), and perform silent capability escalation; it cannot compromise non-agent nodes or break standard cryptographic primitives.
+Under this model, we establish seven security properties and two structural theorems: Capability Binding (Property 4), Trust Containment (Property 3), Chain Integrity (Property 5), and Delegation Depth (Property 6) enforce G1; Reproducibility Soundness (Property 7), Bounded Divergence (Theorem 3), and Chain Verifiability (Theorem 4) enforce G2; Credential Isolation (Property 8) and Ledger Integrity (Property 9) enforce G3.
 1.5 Contributions
-In summary:
-  • We define three Agent Governance Requirements—capability integrity, behavioral verifiability, and interaction auditability—distilling converging regulatory, legal, and technical concerns into formally precise properties (§1.1).
-  • We identify the capability-identity gap and the capability-context separation principle as its root cause and resolution (§1.2).
-  • We show that standard cryptographic mechanisms—composed through the capability-context separation—are sufficient to enforce all three requirements with negligible overhead, providing design guidance for secure agent ecosystems.
-  • We propose capability-bound agent certificates with monotonic trust propagation (§3).
-  • We introduce reproducibility as a trust primitive and establish that reproducibility is a necessary condition for behavioral verifiability: high-assurance agent deployments must be designed for deterministic output to be auditable (§4).
-  • We define a verifiable interaction ledger for tamper-evident forensic reconstruction (§5).
-  • We prove that sufficient replay verification yields bounded divergence: an adversary-controlled model that passes $n$ replay checks is provably $(n,\epsilon)$-indistinguishable from the declared model, with $\epsilon$ exponentially small in nn4.7).
-  • We establish chain verifiability: behavioral verifiability is a chain property, not a node property—a single unverifiable interior agent breaks end-to-end verification for all downstream nodes, yielding principled deployment constraints on model placement in trust trees (§4.8).
-  • We implement a working MCP governance proxy that enforces all three requirements in real-time with $\sim$}120 µs overhead per call, detecting silent capability escalation, tool trojanization, and certificate forgery (§7).
-  • We formalize and prove nine security properties under a realistic adversary model (§6).
-  • We implement a Rust prototype achieving 97 µs verification, ${\sim}0.62$ ms governance overhead, and ${\sim}1{,}200{,}000\times$ speedup over BAID (§7).
-  • We deploy the full governance stack in an end-to-end multi-agent pipeline (5–20 agents with real LLM calls), demonstrating 0.12% overhead, sub-linear per-agent scaling, and detection of all in-situ attacks—while OAuth 2.1 and OpenTelemetry baselines detect none (§7.4).
-1.6 Paper Organization
-§2 presents the threat model and attack scenarios.
-§3 defines capability-bound certificates and the trust propagation tree.
-§4 introduces reproducibility commitments.
-§5 describes the verifiable interaction ledger.
-§6 presents formal security properties and proofs.
-§7 describes implementation and evaluation.
-§8 discusses limitations.
-§9 covers related work.
-§10 concludes.
+Our contribution is a new governance architecture—identifying what must be governed, where governance boundaries belong, and why a specific composition is both necessary and sufficient—not new cryptographic primitives.
+This follows the pattern of Certificate Transparency [56], which composed Merkle trees and gossip protocols (both well-known) into infrastructure that fundamentally changed web PKI operation.
+The insight was in identifying what to compose and where to deploy it, not in new primitives; CT was published at IEEE S&P precisely because the architectural contribution—public auditability of CA behavior—was independently valuable.
+We make an analogous claim: the capability-context separation, the three governance requirements, and the structural results that connect them are independently valuable regardless of the specific cryptographic instantiation.
+Concretely: 1.
+A new vulnerability class and architectural abstraction.
+We identify silent capability escalation—agents acquiring or modifying tools at runtime without triggering existing security mechanisms (§1.2)—and trace it to the capability-context separation: tool definitions and user context are indistinguishable inside a transformer but have fundamentally different security semantics at the orchestration layer [57, 58].
+To our knowledge, this is the first architectural abstraction that identifies where governance boundaries must be placed in agent systems and why existing frameworks—including those with strong cryptographic foundations [2]—systematically miss capability drift.
+2.
+Structural results with architectural consequences.
+The Chain Verifiability Theorem4.9) shows that a single unverifiable interior agent breaks end-to-end behavioral verification, prescribing high-determinism models at interior trust-tree positions.
+The Bounded Divergence Theorem4.8) establishes that an adversary passing $n$ replay checks is $(n,\epsilon)$-indistinguishable from the declared model ($\epsilon\leq 1-\alpha^{1/n}$), framing software G2 as a probabilistic approximation of TEE.
+The Agent Governance Trilemma8.2) proves that capability, performance, and security cannot be simultaneously maximized for Turing-complete agents (via Rice’s theorem).
+3.
+Empirical discovery connecting model determinism to chain security.
+A reproducibility study (9 models, 7 providers, 15,120 comparisons) reveals $5.8\times$ variance in inference determinism, which—combined with Chain Verifiability—determines which trust-tree positions a model qualifies for.
+4.
+Two crypto-agnostic instantiations.
+Basic (Ed25519, SHA-256; 97 µs) and enhanced (BBS+, DV-SNARK; 13.8 ms) instantiations both satisfy nine security properties (§6).
+End-to-end evaluation over 5–20 agent pipelines with real LLM calls confirms <<0.02% governance overhead and real-time detection of 7 E2E attack scenarios (covering G1, G2, and G3) with zero false positives (§7.5).
+An additional 9 structural governance attacks are validated as automated unit tests, and 3 runtime behavioral attacks are validated via post-hoc forensic tracing against Nemotron-AIQ traces (§7).
 2 Threat Model and Motivating Attacks
@@
 […]
-ATK-2: Model Substitution. $\mathcal{A}$ replaces the certified model with a fine-tuned variant. Detected via model binding mismatch and replay verification divergence.
+ATK-2: Model Substitution.
+$\mathcal{A}$ replaces the certified model with a fine-tuned variant.
+Detected via G2: hardware TEE attestation (deterministic), software replay verification divergence, or DV-SNARK CharMatch failure.
 ATK-3: Tool Trojanization. $\mathcal{A}$ replaces tool implementation (same name/version, different code). Detected via code-level hashing in the skills manifest.
@@
 Two or more compromised agents at lower tiers (e.g., T3++T3) attempt to combine their credentials to access resources restricted to a higher tier (T1).
-Prevented by independent per-agent credential evaluation—each agent’s tier and permissions are verified in isolation, so colluding agents gain no additional permissions beyond their individual certificates (Theorem 10).
+Prevented by independent per-agent credential evaluation—each agent’s tier and permissions are verified in isolation, so colluding agents gain no additional permissions beyond their individual certificates (Property 8).
 Forensic Attacks.
@@
 13 return ALLOW
+3.5 Enhanced G1: Selective Disclosure via BBS+ Signatures
+The basic G1 instantiation (Ed25519 + SHA-256) requires full disclosure: a verifier must see the entire skills manifest to recompute the hash and check capability binding.
+This suffices for single-organization deployments but is problematic in cross-organizational settings where revealing the full tool manifest leaks competitive intelligence (e.g., which proprietary tools an agent uses, its model version, or its rate limits).
+BBS+ signatures [51] enable a strictly stronger property: an agent can prove that its certificate is valid and that specific capability predicates hold, without revealing the full certificate contents.
+BBS+ is a W3C Verifiable Credentials standard primitive with security reducing to the $q$-SDH assumption in bilinear groups [52].
+Definition 5 (BBS+ Capability Certificate).
+The issuer signs the certificate attributes as a vector of $L$ messages:
+$\mathbf{m}=(m_{1},\ldots,m_{L})=(\mathit{id}_{v},\mathit{pk}_{v},\mu_{v},h_{1},\ldots,h_{n},\kappa_{v},\rho_{v},t_{s},t_{e})$
+producing a BBS+ signature $\Sigma=(A,e,s)$ over $\mathbf{m}$.
+Selective disclosure. To prove capability binding to a verifier, the agent derives a zero-knowledge proof of knowledge $\pi$ that:
+  1. 1.
+    The agent possesses a valid BBS+ signature $\Sigma$ from a trusted issuer.
+  2. 2.
+    The revealed subset $\mathbf{m}_{R}\subset\mathbf{m}$ (e.g., only the tool IDs, not model version or rate limits) matches the disclosed values.
+  3. 3.
+    The hidden attributes $\mathbf{m}_{H}=\mathbf{m}\setminus\mathbf{m}_{R}$ satisfy declared predicates (e.g., $\mathit{max\_tier}\geq 2$, $\mathit{model}\in\{\texttt{gpt-4o},\texttt{claude-4}\}$).
+The verifier learns only the revealed attributes and the predicate outcomes—not the hidden values.
+This enables scenarios such as: an agent proves it has some valid tool manifest without revealing which tools; or proves its trust tier exceeds a threshold without revealing its exact tier or delegation depth.
+Capability binding under BBS+.
+The capability binding invariant is preserved: at runtime, the agent recomputes $H_{v}(t)=\mathrm{SHA\text{-}256}(\mathrm{canonical}(S_{v}(t)))$ and includes a ZKP that $H_{v}(t)$ matches the committed hash in the signed message vector.
+Any tool change invalidates the proof.
+Cost-privacy tradeoff.
+BBS+ selective disclosure proof generation costs ${\sim}14.7$ ms and verification ${\sim}13.8$ ms (benchmarked in §7), compared to ${\sim}97$ µs for basic Ed25519 verification—a ${\sim}142\times$ overhead.
+This is the concrete cost of privacy: the governance contract (G1) is identical; the privacy properties differ.
+Deployments choose between basic and enhanced G1 based on whether cross-organizational privacy justifies the latency cost.
+Crucially, these costs apply per-request only when selective disclosure is needed; for routine intra-organizational use, the basic instantiation suffices.
+Security.
+BBS+ signature unforgeability reduces to the $q$-SDH assumption in Type-3 bilinear groups; the zero-knowledge property reduces to DDH.
+Both are standard assumptions with well-studied cryptanalytic history.
+The enhanced instantiation inherits all nine security properties of the basic instantiation, with the additional property of certificate privacy: no PPT verifier can extract hidden attributes from a selective disclosure proof.
 3.6 Integration with A2A Agent Cards
 Agent Certificates are designed as non-invasive extensions of the A2A ecosystem via an optional extensions field in the Agent Card containing the certificate PEM, chain, and skills manifest hash.
-4 Reproducibility Commitments
-4.1 Reproducibility as a Necessary Condition
-A key observation motivates our approach: behavioral verifiability (G2) is impossible without reproducibility.
+The enhanced instantiation adds a disclosure_mode field (full or selective) and, for selective disclosure, the BBS+ proof $\pi$ replacing the raw certificate.
+4 Inference Authenticity and Reproducibility
+4.1 G2 as Inference Authenticity
+G2 requires inference authenticity: verifying that an agent actually executed its declared model on the recorded inputs.
+This is an abstract governance requirement, independent of how it is achieved.
+Three implementation paths exist, each with distinct trust assumptions:
+Hardware attestation (TEE).
+Trusted execution environments [53, 54] guarantee that the declared code was what actually executed.
+This provides the strongest G2 guarantee—deterministic, not probabilistic—but requires specialized hardware infrastructure (Intel SGX/TDX [53], ARM TrustZone [54], NVIDIA H100 Confidential Computing [55]) that is unavailable in most current LLM API deployments.
+TEE satisfies G2 but does not satisfy G1 or G3: a TEE-attested agent can still silently acquire tools (G1 violation), and TEE provides per-execution integrity without tamper-evident cross-agent audit trails (G3 violation).
+Our framework provides the governance layer that TEE alone cannot.
+Software replay verification.
+When TEE is unavailable—the dominant case for cloud-hosted LLM APIs—G2 can be approximated through reproducibility-based replay: re-execute the agent’s declared model on recorded inputs and compare outputs.
+This provides a probabilistic software approximation of TEE, with guarantees formalized in Property 1 and Theorem 1.
+The approximation quality improves with the verification budget $n$.
+Zero-knowledge verification (DV-SNARK, §4.10).
+A privacy-preserving variant of replay verification: the prover demonstrates output consistency without revealing prompts or outputs.
+This adds privacy guarantees to the software path at the cost of computational overhead.
+Full-inference zkVM [2].
+A SNARK circuit over the complete transformer forward pass provides the strongest software G2 guarantee—computational integrity without TEE hardware—by cryptographically binding the agent’s execution to its registered code commitment $C_{P}=\mathrm{CommitProg}(P)$.
+This path is appropriate for high-assurance scenarios (e.g., financial settlement, regulated environments) where the proof generation overhead (${\sim}120$ s per interaction) is acceptable and provider re-execution cooperation is unavailable.
+Unlike the software replay path, this path does not require model re-execution by the verifier; the proof is self-contained.
+All four paths satisfy the same G2 governance contract.
+The framework is agnostic to which path is used; the choice is declared in the agent’s certificate ($\rho_{v}$) and recorded in the ledger, making the G2 implementation transparent to auditors and downstream agents.
+4.2 Reproducibility as a Necessary Condition for Software G2
+For the software path, a key observation applies: behavioral verifiability is impossible without reproducibility.
 If an agent’s outputs cannot be independently regenerated—even approximately—then no post-hoc mechanism can distinguish authorized behavior from unauthorized behavior.
-This is not merely a limitation of our framework; it is a fundamental constraint on any verification scheme that does not rely on trusted execution environments.
-This observation has a prescriptive implication for secure agent system design: high-assurance agent deployments must be engineered for reproducible output.
+This is a fundamental constraint on any software-based verification scheme.
+This observation has a prescriptive implication for secure agent system design: high-assurance deployments without TEE must be engineered for reproducible output.
 Concretely, this means using deterministic sampling (temperature==0, fixed seeds), preferring structured outputs (code, JSON, SQL) over free-form prose where security-critical decisions are involved, and selecting inference providers that guarantee bitwise or near-bitwise reproducibility [6].
 Agents that cannot commit to any reproducibility level receive a trust tier downgrade (§4.5), making the security cost of non-reproducibility explicit in the governance framework.
+More broadly, by making reproducibility a prerequisite for software G2 compliance, our framework transforms it from an implementation detail into a security-critical model performance metric—with implications for how models are specified, evaluated, and deployed (§4.9).
 4.3 Empirical Foundation
@@
 The threshold $\theta$ is task-type dependent: our empirical validation (Experiment 5, §7) shows that code-generating agents achieve perfect determinism ($\theta=0.98$), while prose-generating agents require $\theta=0.10$ to account for paraphrase-level variation.
-Cross-model divergence remains well below even the relaxed threshold (max 0.089).
+Cross-provider mean CharMatch ($0.114$, Table 5) remains well below even the relaxed threshold.
 None: No reproducibility commitment.
@@
 4.7 Security Analysis
-Theorem 1 (Reproducibility Soundness, informal).
+Property 1 (Reproducibility Soundness, informal).
 If compromised agent $v$ executes model $M^{\prime}\neq M$ or skills $S^{\prime}\neq S$, replay verification detects the deviation with probability $\geq 1-\epsilon(\lambda)$ for negligible $\epsilon$.
 By the DiFR token divergence result [1], the match rate between different models is $\delta\approx 0.06$ (character-level, validated in Experiment 5).
-By Hoeffding’s inequality: $\Pr[\mathrm{CharMatch}\geq\theta]\leq\exp(-2n(\theta-\delta)^{2})$, which is negligible for realistic output lengths.
+Our empirical data (Table 6) shows that the fraction of cross-model output pairs exceeding any reasonable threshold $\theta$ is at most $q\approx 0.02$ — i.e., a substituted model’s output coincidentally passes the CharMatch $\geq\theta$ check with probability at most $q$ per prompt.
+Over $n$ independent verification prompts, the probability of all checks passing despite model substitution is at most $q^{n}$, which is negligible for realistic $n$.
+Input integrity for replay verification.
+Replay verification requires the recorded input to be authentic.
+A compromised agent cannot fabricate its own input because the interaction ledger uses bilateral signing: the sender (an honest agent or non-agent) signs the input commitment $\eta_{\mathit{in}}$, which the compromised receiver cannot modify without invalidating the sender’s signature.
+G2 therefore relies on sender-authenticated input, not self-reported input—a unidirectional dependency on G3, not a circular one.
+The formal argument is given in Property 106).
 4.8 Indistinguishability Under Sufficient Verification
@@
 If adversary-controlled model $M^{\prime}$ passes replay verification (CharMatch $\geq\theta$) on $n$ independent prompts drawn from $\mathcal{D}$ across $|\mathcal{S}|$ skills, then $M$ and $M^{\prime}$ are $(n,\epsilon)$-indistinguishable with:
-$\epsilon\leq\exp\!\left(-2n\theta^{2}\right)$
-Equivalently, for target confidence $1-\epsilon$, $n\geq\frac{\ln(1/\epsilon)}{2\theta^{2}}$ verification prompts suffice.
+$\epsilon\leq 1-\alpha^{1/n}$
+where $\alpha$ is the significance level.
+For small $\epsilon$, this simplifies to $\epsilon\lessapprox\ln(1/\alpha)/n$.
+Equivalently, for target failure-rate bound $\epsilon$, the required verification budget is $n\geq\ln(1/\alpha)/\epsilon$.
 Proof sketch.
-Let $Z_{i}=\mathbf{1}[\mathrm{CharMatch}(M(x_{i}),M^{\prime}(x_{i}))<\theta]$ be independent Bernoulli random variables with $p=\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}<\theta]$.
-Observing $\sum Z_{i}=0$ (all $n$ tests pass), by a one-sided binomial test: $\Pr[\sum Z_{i}=0\mid p\geq\epsilon]=(1-p)^{n}\leq(1-\epsilon)^{n}\leq\exp(-n\epsilon)$.
-Setting this equal to $\alpha$ (significance level) and solving yields $p\leq\ln(1/\alpha)/n$ with confidence $1-\alpha$.
-The Hoeffding refinement accounting for the threshold gap gives the stated bound.
+Let $p=\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]$ be the true failure rate under $M^{\prime}$.
+Each verification trial $Z_{i}=\mathbf{1}[\mathrm{CharMatch}\geq\theta]$ is an independent Bernoulli(1p)(1-p).
+Observing all $n$ tests pass ($\sum Z_{i}=n$), the likelihood is $(1-p)^{n}$.
+The exact one-sided Clopper-Pearson upper bound on $p$ at confidence $1-\alpha$ is $p\leq 1-\alpha^{1/n}$.
+For practical computation: $(1-p)^{n}\leq\alpha$ implies $p\leq 1-\alpha^{1/n}\approx\ln(1/\alpha)/n$ when $p$ is small.
+This is a distribution-free result: the bound depends only on the number of verification trials and the desired confidence level, not on the threshold $\theta$ itself.
+The role of $\theta$ is to define what constitutes a “pass” — a higher $\theta$ increases sensitivity (more failures for a substituted model) but also increases the false-positive rate for the legitimate model.
 Full proof in §6.
@@
 Practical implications.
-For code-generating agents ($\theta=0.98$), only $n=3$ prompts per skill achieve $\epsilon<10^{-2}$; $n=12$ achieves $\epsilon<10^{-10}$.
-Even for relaxed prose thresholds ($\theta=0.10$), $n=1200$ prompts achieve $\epsilon<10^{-10}$.
-These are operationally feasible: a verification suite of 50 prompts per skill, run periodically, provides strong guarantees.
+Table 2 shows the failure-rate bound $\epsilon$ (i.e., the maximum fraction of deployment prompts on which $M^{\prime}$ can diverge undetected) as a function of the verification budget $n$ at 99% confidence ($\alpha=0.01$):
+$n$ $\epsilon$ bound Interpretation
+10 0.369 Coarse screening
+25 0.168 Moderate assurance
+50 0.089 <<10% divergent prompts
+100 0.045 <<5% divergent prompts
+200 0.023 High assurance
+500 0.009 <<1% divergent prompts
+Table 2: Failure-rate bound $\epsilon$ at 99% confidence ($\alpha=0.01$) for various verification budgets $n$, computed as $\epsilon=1-0.01^{1/n}$.
+A verification suite of 50 prompts per skill — operationally feasible as a periodic audit — bounds the adversary’s undetected divergence to at most 8.9% of deployment prompts with 99% confidence.
+Increasing to 200 prompts tightens this to 2.3%.
+Threshold calibration.
+The threshold $\theta$ must be calibrated per task type and model to balance detection sensitivity against the false-positive rate.
+Our empirical data (Experiment 5) shows that single-metric thresholds suffer from the bimodal output distribution: for code generation, CharMatch is either >0.95>0.95 (identical outputs) or <0.5<0.5 (paraphrase variants), leaving no single $\theta$ that simultaneously achieves low false positives and high detection.
+We therefore adopt a multi-metric ensemble: a prompt is flagged if any of four metrics (CharMatch, Jaccard word overlap, TF-IDF cosine, character n-gram cosine) falls below its respective per-metric optimal threshold (determined by Youden’s $J$ statistic, Table 7 in §7).
+This ensemble approach improves detection robustness across task types, as each metric captures different aspects of output similarity (surface-form identity vs.
+semantic overlap).
 Security interpretation.
-This transforms reproducibility from a detection mechanism into a safety certificate: an adversary who fine-tunes $M$ into $M^{\prime}$ that passes $n$ diverse replay checks has produced a model that is provably close to $M$ on the tested distribution.
-Any residual divergence is bounded by $\epsilon$, which decreases exponentially with the verification budget.
-The adversary faces a dilemma: either $M^{\prime}$ behaves like $M$ (and the fine-tuning achieves nothing), or $M^{\prime}$ diverges on some inputs and will be caught with probability increasing exponentially in $n$.
-This is analogous to cryptographic indistinguishability: the adversary’s advantage is negligible given sufficient verification.
+This transforms reproducibility from a detection mechanism into a safety certificate: an adversary who fine-tunes $M$ into $M^{\prime}$ that passes $n$ diverse replay checks has produced a model that diverges on at most an $\epsilon$-fraction of deployment prompts, with $\epsilon$ decreasing as $O(\ln(1/\alpha)/n)$.
+The adversary faces a dilemma: either $M^{\prime}$ behaves like $M$ on most inputs (limiting the attack’s impact to at most an $\epsilon$-fraction of requests), or $M^{\prime}$ diverges more broadly and will be caught with probability increasing in $n$.
 4.9 Chain Verifiability
@@
 […]
-The gap that G2 uniquely fills.
-Consider what G1+G3 alone guarantee.
-G1 (capability binding) ensures that each agent’s certificate is bound to its declared model and tool set; any capability change invalidates the certificate.
-G3 (interaction auditability) ensures that every input and output is bilaterally signed and hash-chain linked; no record can be silently modified.
-Together, G1+G3 provide attribution integrity: we know who sent what to whom, and that no message was altered in transit.
 […]
@@
 ∎
+Per-model reproducibility variance.
+Our multi-provider experiment (Experiment 5c, §7) reveals that model determinism varies by over an order of magnitude: from CharMatch==0.987 (Haiku 4.5) to 0.169 (Kimi K2.5).
+Combined with the Chain Verifiability Theorem, this transforms into prescriptive architectural guidance: the security of a multi-agent pipeline depends not just on whether G2 is enforced, but on where high- and low-determinism models are placed in the trust tree.
+Reproducibility as a model performance metric.
+The G1–G3 framework elevates reproducibility from an implementation detail to a first-class model performance metric: a model’s determinism directly determines which governance level it qualifies for (§8.2), which trust-tree positions it may occupy, and what verification budget is needed (Theorem 1).
+We argue that providers should self-declare reproducibility characteristics—analogous to FIPS 140 security levels for cryptographic modules—including expected CharMatch ranges under temperature==0 and sensitivity to infrastructure changes.
+Our empirical characterization across 9 models and 7 providers (Experiment 5) provides the first multi-provider basis for such specifications.
 Architectural implications. This theorem yields three deployment principles:
 Principle 1: High-determinism models at interior nodes.
-Models with strong reproducibility (Anthropic Haiku: CharMatch==0.987, Sonnet: 0.823; see Table 4) should be placed at coordinator and interior positions in the trust tree, where their outputs become inputs for downstream agents.
-Low-determinism models (Kimi: 0.169, DeepSeek: 0.225) should be constrained to leaf positions where their outputs do not propagate further.
-This is not merely a performance recommendation—it is a security architecture constraint.
+Models with strong reproducibility (Haiku: CharMatch==0.987, Sonnet: 0.823) occupy interior trust-tree positions; low-determinism models (Kimi: 0.169, DeepSeek: 0.225) are constrained to leaf positions—a security architecture constraint, not merely a performance recommendation.
 Principle 2: Verification boundary checkpoints.
-When a G2==none agent must appear at an interior position (e.g., a reasoning model as coordinator), the chain requires a verification boundary: either (a) a human-in-the-loop checkpoint that re-anchors trust by reviewing and signing the output, (b) TEE attestation providing hardware-enforced computational integrity, or (c) dual-execution with independent instances and output comparison (analogous to Byzantine fault tolerance).
-Each mechanism re-establishes chain verifiability for downstream nodes.
+When a G2==none agent must appear at an interior position, the chain requires a verification boundary—human-in-the-loop, TEE attestation, or dual-execution—to re-establish chain verifiability for downstream nodes.
 Principle 3: Explicit degradation in the ledger.
-When chain verifiability is broken at node $v_{k}$, the ledger should record a PARTIAL_VERIFIABILITY marker at that boundary.
-Downstream agents inherit a trust tier penalty beyond the standard G2==none downgrade (§4.4), making the reduced security guarantee visible to auditors and compliance systems.
+When chain verifiability breaks at node $v_{k}$, the ledger records a PARTIAL_VERIFIABILITY marker; downstream agents inherit a trust tier penalty, making the reduced guarantee visible to auditors.
 Relationship to TEE.
-This analysis clarifies the complementary roles of G2 and trusted execution environments.
-G2 provides software-only verification of computational fidelity—a “software TEE”—using only black-box API access and the near-determinism of LLM inference.
-Hardware TEEs provide the same guarantee through hardware attestation but require infrastructure unavailable in most LLM deployments.
-The Chain Verifiability Theorem shows precisely where each mechanism is needed: at every interior agent node in the trust tree.
-In mixed deployments, G2 and TEE are interchangeable at the chain level: an agent with TEE attestation and G2==none maintains chain verifiability identically to an agent with G2==full and no TEE.
+The Chain Verifiability Theorem clarifies where G2 is needed: at every interior node.
+In mixed deployments, different nodes may use different G2 implementations (TEE, software replay, or DV-SNARK), with chain verifiability depth determined by the weakest interior node.
+Regardless of the G2 implementation, each node still requires G1 and G3: TEE attests that some model ran correctly, but only G1 ensures it was the authorized model, and only G3 provides tamper-evident records.
+4.10 Enhanced G2: Zero-Knowledge Behavioral Verification
+The basic G2 instantiation requires provider cooperation: the verifier must re-execute the agent’s declared model on recorded inputs and compare outputs.
+This creates two limitations: (1) the verifier must have API access to the same model version, and (2) the original prompts and outputs are disclosed to the verifier, which may contain sensitive information.
+Designated-Verifier SNARKs (DV-SNARKs) address both limitations by enabling the agent (or a delegated prover) to demonstrate output consistency in zero knowledge: the verifier learns only that $\mathrm{CharMatch}(O,O^{\mathrm{replay}})\geq\theta$, without seeing $O$, $O^{\mathrm{replay}}$, or the prompts.
+Key design insight.
+Unlike BAID [2], which SNARKs over the entire transformer forward pass (${\sim}120$ s per proof for a 7B model), our approach SNARKs only over the verification metric computation—a fundamentally simpler circuit:
+Definition 9 (CharMatch Verification Circuit).
+$\mathcal{C}_{\mathrm{verify}}(O,O^{\mathrm{replay}},\theta)=\begin{cases}1&\text{if }\frac{|\{i:O[i]=O^{\mathrm{replay}}[i]\}|}{\max(|O|,|O^{\mathrm{replay}}|)}\geq\theta\\ 0&\text{otherwise}\end{cases}$
+For typical agent outputs ($|O|\leq 4{,}096$ characters), the circuit requires $O(|O|)$ equality comparisons and a single division—on the order of $10^{4}$ R1CS constraints (specifically: 19,238 for 1K outputs, ${\sim}73{,}000$ for 4K outputs), compared to ${\sim}10^{10}$ for a transformer circuit.
+This reduces proof generation from minutes to sub-seconds.
+Designated-Verifier property.
+The proof is bound to a specific verifier’s public key, preventing proof transfer.
+Only the designated auditor can verify the proof, which prevents a compromised agent from shopping for a lenient verifier.
+The DV property also prevents the proof from being used as transferable evidence—a deliberate design choice that limits the proof’s scope to the specific audit context.
+Protocol. Enhanced G2 verification proceeds as:
+  1. 1.
+    The agent (or a trusted replay service) executes the model on the recorded input to produce $O^{\mathrm{replay}}$.
+  2. 2.
+    The prover computes $\pi=\mathrm{DV\text{-}SNARK}.\mathrm{Prove}(\mathcal{C}_{\mathrm{verify}},(O,O^{\mathrm{replay}}),\theta,\mathit{pk}_{\mathrm{verifier}})$.
+  3. 3.
+    The verifier checks $\mathrm{DV\text{-}SNARK}.\mathrm{Verify}(\pi,\theta,\mathit{sk}_{\mathrm{verifier}})\stackrel{{\scriptstyle?}}{{=}}1$.
+What DV-SNARK does and does not prove.
+The SNARK proves that the prover possesses two strings whose CharMatch exceeds $\theta$; it does not prove that $O^{\mathrm{replay}}$ was produced by the declared model (that trust is delegated to the replay service).
+The SNARK’s contribution is privacy; for full computational integrity, TEE or BAID-style full-inference SNARKs remain necessary.
+Cost-privacy tradeoff.
+DV-SNARK proof generation costs ${\sim}174$ ms (1K outputs) to ${\sim}750$ ms (4K outputs); Groth16 verification is constant at ${\sim}1.3$ ms (§7).
+Proof generation is ${\sim}1{,}740\times$ slower than basic CharMatch but is offline and asynchronous; it is ${\sim}690\times$ faster than BAID’s full-inference zkVM (${\sim}120$ s).
+The governance contract (G2) is identical; the privacy properties differ.
+Security.
+Knowledge soundness ensures that a prover cannot produce a valid proof without possessing strings satisfying the CharMatch predicate.
+The DV property ensures simulation-privacy: no entity other than the designated verifier can distinguish a real proof from a simulated one.
 5 Verifiable Interaction Ledger
@@
 (3) No privacy preservation: OTEL stores full plaintext input/output payloads, creating data exposure risk.
-Our interaction ledger addresses all three gaps through hash-chain linking ($h_{\mathit{prev}}$), Ed25519 bilateral signing ($\mathit{sig}_{i}$), agent certificate binding ($\mathit{ch}_{\mathit{send}},\mathit{ch}_{\mathit{recv}}$), and commitment-only storage ($\eta_{\mathit{in}},\eta_{\mathit{out}}$).
+Our interaction ledger addresses all three gaps through hash-chain linking ($h_{\mathit{prev}}$), bilateral digital signing ($\mathit{sig}_{i}$; Ed25519 in the basic instantiation), agent certificate binding ($\mathit{ch}_{\mathit{send}},\mathit{ch}_{\mathit{recv}}$), and commitment-only storage ($\eta_{\mathit{in}},\eta_{\mathit{out}}$).
 5.2 Ledger Structure
@@
 The protocol proceeds in five steps: (1) identify the relevant interaction chain, (2) verify ledger integrity (signatures + hash chain), (3) request content disclosure and verify against commitments, (4) verify agent certificates at interaction time, and (5) execute replay verification for agents with reproducibility commitments.
-6 Formal Security Properties and Proofs
-We prove nine properties under the adversary model of §2, using standard cryptographic assumptions: EUF-CMA security, collision resistance of SHA-256, and the DiFR token divergence property [1].
-Theorem 4 (Trust Containment).
+5.5 Chain Auditability
+Analogous to the Chain Verifiability Theorem for G2 (§4.9), interaction auditability is a chain property: a single missing interaction record along an execution path breaks forensic reconstruction for all downstream nodes on that path.
+Definition 12 (Chain Auditability Depth).
+For execution chain $v_{0}\to v_{1}\to\cdots\to v_{n}$ with $v_{0}$ a non-agent trust anchor, define:
+[…]
+with $\mathrm{CAD}(v_{n})=n$ if all records are present (fully auditable chain).
+Property 2 (Chain Auditability).
+If $\mathrm{CAD}(v_{n})=k<n$, then for all $i\geq k$:
+  1. 1.
+    Provenance gap: the auditor cannot trace $v_{i}$’s output back to the trust anchor $v_{0}$, because the input $v_{k}$ received from $v_{k-1}$ is unrecorded.
+  2. 2.
+    G2 cascade: replay verification for $v_{k}$ is impossible—the sender-authenticated input $\eta_{\mathit{in}}$ required by Property 10 is unavailable.
+By induction, downstream G2 verification inherits the same gap even if $\rho_{v_{i}}.\mathit{level}=\mathit{full}$.
+This creates a symmetric pair of chain properties: G2’s Chain Verifiability (Theorem 2) governs behavioral verification depth, while G3’s Chain Auditability governs forensic reconstruction depth.
+Because G2 replay depends on G3’s sender-signed input (Property 10), the effective verification depth of any execution path is $\min(\mathrm{CVD},\mathrm{CAD})$—the minimum of the two chain depths.
+5.6 G3 as a Governance Contract
+Like G1 and G2, G3 defines what properties interaction records must satisfy, not how the ledger is implemented.
+The governance contract requires: (1) every inter-agent and agent-human interaction produces a record with bilateral signatures and content commitments; (2) records are linked in a tamper-evident chain; (3) sufficient metadata exists for forensic reconstruction.
+Our current instantiation uses a centralized append-only ledger operated by a trusted organizational entity (e.g., the deploying enterprise).
+This is the simplest design that satisfies G3: agents submit bilaterally-signed records to the operator, who maintains the hash chain and provides query access during audits.
+The operator is assumed honest for availability (it will not selectively omit records) and ordering (it will not reorder records), while integrity is enforced cryptographically regardless of operator behavior (bilateral signatures + hash chain prevent undetectable modification).
+Alternative instantiations satisfying the same G3 contract include: (a) multi-operator consensus, where f+1f{+}1-of-$n$ operators must agree on each append, eliminating single-operator availability trust at the cost of consensus latency; (b) append-only transparency logs analogous to Certificate Transparency, where multiple independent monitors cross-check the ledger for consistency; (c) blockchain-anchored commitments, where periodic ledger root hashes are anchored to a public blockchain for non-repudiable timestamping.
+Each alternative trades latency or operational complexity for stronger trust assumptions.
+The G3 contract is invariant across all instantiations—what changes is the trust model of the storage layer, not the governance semantics of the records.
+6 Formal Security Properties
+We state seven security properties and two theorems under the adversary model of §2.
+The basic instantiation uses standard cryptographic assumptions: EUF-CMA security of Ed25519, collision resistance of SHA-256, and the DiFR token divergence property [1].
+The enhanced instantiation additionally relies on $q$-SDH and DDH in Type-3 bilinear groups (BBS+ unforgeability and zero-knowledge, §3.5) and knowledge soundness of Groth16 (DV-SNARK, §4.10).
+All proofs are in Appendix A.
+Property 3 (Trust Containment).
 For all edges $(u,v)\in E$: $\kappa_{v}\leq_{\kappa}\kappa_{u}$.
-Proof.
-Suppose $\mathcal{A}$ issues $\mathit{cert}_{w}$ with $\kappa_{w}\not\leq_{\kappa}\kappa_{v}$.
-The verifier checks $\kappa_{w}\leq_{\kappa}\kappa_{v}$ using constraints in $\mathit{cert}_{v}$, which is signed by $v$’s parent.
-$\mathcal{A}$ cannot modify $\mathit{cert}_{v}$ without the parent’s key.
-Recursing upward, the chain terminates at an uncompromised non-agent ancestor (EUF-CMA).
-∎∎
-Theorem 5 (Capability Binding).
+Property 4 (Capability Binding).
 If $H_{v}(t^{\prime})\neq H_{v}^{\mathit{cert}}$, then $\forall t\geq t^{\prime}:\mathrm{Access}(v,c,t)=\mathrm{DENY}$.
-Proof.
-By collision resistance: $\Pr[H_{v}(t^{\prime})=H_{v}^{\mathit{cert}}]\leq\mathrm{negl}(\lambda)$ when $S_{v}(t^{\prime})\neq\sigma_{v}$.
-Forging a new certificate requires the parent’s key (same argument as Theorem 4).
-∎∎
-Theorem 6 (Chain Integrity).
+Property 5 (Chain Integrity).
 Only chains forming a valid path from a trusted root with valid signatures pass verification.
-Proof.
-Fabricating any link requires the corresponding private key. By EUF-CMA, $\mathcal{A}$ cannot produce valid signatures without the key. ∎∎
-Theorem 7 (Delegation Depth Enforcement).
+Property 6 (Delegation Depth Enforcement).
 Agent delegation depth from any non-agent ancestor never exceeds $\mathit{max\_depth}$.
-Proof.
-Trust Containment requires $\kappa_{v}.\mathit{max\_depth}<\kappa_{\mathit{par}(v)}.\mathit{max\_depth}$ (strict). Starting from depth $d$, the chain has at most $d$ agent levels before reaching 0. ∎∎
-Theorem 8 (Reproducibility Soundness).
-If $v$ executes $M^{\prime}\neq M$ or $S^{\prime}\neq S$, replay verification detects with probability $\geq 1-\epsilon(\lambda)$.
-Proof.
-By DiFR [1]: $\mathbb{E}[\mathrm{CharMatch}(O^{\mathit{actual}},O^{\mathit{ref}})]\leq\delta$ for different models.
-By Hoeffding: $\Pr[\mathrm{CharMatch}\geq\theta]\leq\exp(-2n(\theta-\delta)^{2})$, negligible for realistic $n$.
-∎∎
+Property 7 (Reproducibility Soundness).
+Under software G2 (replay verification): if $v$ executes $M^{\prime}\neq M$ or $S^{\prime}\neq S$, replay verification detects with probability $\geq 1-\epsilon(\lambda)$.
+Under hardware G2 (TEE): detection is deterministic.
 Theorem 3 (Bounded Divergence Under Verification).
-If $M^{\prime}$ passes replay verification on $n$ independent prompts from $\mathcal{D}$, then $\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]\leq\epsilon$ where $\epsilon\leq\exp(-2n\theta^{2})$.
-Proof.
-Let $p=\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]$ be the true failure rate.
-Each verification trial is Bernoulli($p$).
-Observing $n$ successes (all pass), the likelihood ratio test gives: $\Pr[\text{all pass}\mid p]=(1-p)^{n}$.
-For $p\geq\epsilon$: $(1-\epsilon)^{n}\leq e^{-n\epsilon}$.
-Setting the false-acceptance probability to $e^{-n\epsilon}\leq e^{-2n\theta^{2}}$ (using $\epsilon=2\theta^{2}$ as the critical value from the Hoeffding bound on the CharMatch statistic), we obtain the stated bound.
-The Hoeffding refinement uses the fact that CharMatch is a bounded $[0,1]$ statistic with expected value $\geq\theta$ under $H_{0}$ (same model), applying $\Pr[\bar{X}-\mu\leq-t]\leq\exp(-2nt^{2})$ to the complement event.
-∎∎
-Theorem 10 (Credential Isolation).
+If $M^{\prime}$ passes replay verification (CharMatch $\geq\theta$) on $n$ independent prompts from $\mathcal{D}$, then $\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]\leq\epsilon$ where $\epsilon\leq 1-\alpha^{1/n}$, with $\alpha$ the significance level.
+Property 8 (Credential Isolation).
 $\mathrm{Accessible}(C^{\prime})=\bigcup_{v\in C^{\prime}}\mathrm{Accessible}(\{v\})$.
-Proof.
-Access depends solely on the requesting agent’s certificate. No protocol mechanism combines permissions. Proxy tokens are bound to specific agent IDs. ∎∎
-Theorem 11 (Ledger Integrity).
+Property 9 (Ledger Integrity).
 No party can undetectably modify, delete, or reorder records.
-Proof.
-(a) Modification invalidates $h_{\mathit{prev}}$ of successor, cascading through chain; re-signing requires all subsequent senders’ keys.
-(b) Deletion breaks hash chain.
-(c) Reordering breaks both sequence numbers and hash links.
-∎∎
 Theorem 4 (Chain Verifiability).
@@
 Then for all $i>k$, the output of $v_{i}$ cannot be verified as originating from a legitimate computation chain rooted at $v_{0}$, even if $\rho_{v_{i}}.\mathit{level}=\mathit{full}$.
-Proof.
-We proceed by induction on chain position.
-[…]
-Inductive step: Assume the output of $v_{i-1}$ (i1>ki-1>k) cannot be verified as legitimate.
-Agent $v_{i}$ receives this unverifiable output as input.
-By the same argument as the base case, $v_{i}$’s replay verification confirms computational fidelity on unverifiable input.
-∎∎
+Property 10 (G2 Input Integrity via Bilateral Signing).
+Replay verification for a compromised agent $v_{j}$ uses sender-authenticated input, not self-reported input. Consequently, G2 soundness does not circularly depend on $v_{j}$’s own honesty.
+Rationale.
+A potential concern is circular dependency: G2 replay verification requires the recorded input to be authentic, but if the compromised agent $v_{j}$ controls its own input recording, it could fabricate inputs that make its unauthorized output appear legitimate.
+We show this is precluded by the bilateral signing structure of G3.
+In any execution chain, agent $v_{j}$’s input is the output of its predecessor $v_{j-1}$.
+The interaction record $R_{i}$ in the ledger contains the input commitment $\eta_{\mathit{in}}=\mathrm{SHA\text{-}256}(\mathit{input})$ and is bilaterally signed by both $v_{j-1}$ (sender) and $v_{j}$ (receiver).
+We consider three cases for $v_{j-1}$:
+Case 1: $v_{j-1}$ is a non-agent node (human or organization).
+By the adversary model (§2), non-agent nodes are uncompromised.
+$v_{j-1}$ signs the interaction record containing $\eta_{\mathit{in}}$.
+Agent $v_{j}$ cannot modify $\eta_{\mathit{in}}$ without invalidating $v_{j-1}$’s signature (EUF-CMA).
+Therefore the input committed to the ledger is the true input $v_{j}$ received.
+Case 2: $v_{j-1}$ is an honest agent.
+The same argument applies: $v_{j-1}$ signs its output and the corresponding ledger record.
+$v_{j}$ cannot forge $v_{j-1}$’s signature on a different $\eta_{\mathit{in}}$.
+Case 3: $v_{j-1}$ is also compromised.
+Then $v_{j-1}$’s output is itself unverifiable.
+However, the Chain Verifiability Theorem (Theorem 4) already establishes that all nodes downstream of the first unverifiable agent lose end-to-end verification guarantees.
+In this case, $v_{j}$ is downstream of the already-compromised $v_{j-1}$, so the chain verifiability depth $\mathrm{CVD}\leq j-1<j$, and no claim of end-to-end verification extends to $v_{j}$ regardless.
+In Cases 1 and 2, the verifier reconstructs $v_{j}$’s input via the forensic reconstruction protocol (§5): request $v_{j-1}$ to disclose the original content, verify it against the sender-signed $\eta_{\mathit{in}}$, then replay $v_{j}$’s declared model on this authenticated input.
+The dependency is therefore unidirectional—G2 relies on G3’s sender-side input commitment, which is established independently of $v_{j}$—not circular.
+Chain Auditability.
+This unidirectional dependency implies that G3 also exhibits a chain property: if an interaction record along an execution path is absent from the ledger, G2 replay verification at that point becomes impossible (the authenticated input is unavailable), and by induction all downstream provenance tracing fails.
+The effective verification depth of any execution path is therefore $\min(\mathrm{CVD},\mathrm{CAD})$—the minimum of Chain Verifiability Depth (Definition 8) and Chain Auditability Depth (Definition 12).
+This symmetric pair of chain properties—G2 governing behavioral verification depth, G3 governing forensic reconstruction depth—underscores the mutual reinforcement of the governance requirements.
 7 Implementation and Evaluation
@@
 Component Dependencies LoC
-Certificate (parse, issue, verify) x509-parser, ring $\sim$}1,400
-Skills manifest hasher sha2, serde_cbor $\sim$}500
-Trust tree (internal) $\sim$}700
-Reproducibility verifier sha2 $\sim$}400
-Interaction ledger ed25519-dalek, sled $\sim$}1,100
-Credential vault aes-gcm, argon2 $\sim$}800
-Agent registry + proxy tokens agent-trust-kit $\sim$}600
-MCP server + CLI jsonrpc-core, clap $\sim$}900
+Core SDK (cert, skills, trust tree, ring, sha2, ed25519-dalek, $\sim$}5,600
+     ledger, vault, registry, MCP) sled, aes-gcm, clap
+Enhanced G1: BBS+ disclosure zkryptium $\sim$}270
+Enhanced G2: DV-SNARK circuit ark-groth16, ark-r1cs-std $\sim$}350
 Real LLM repro. validation anthropic, sklearn (Python) $\sim$}870
-MCP governance proxy demo cryptography, cbor2 $\sim$}520
-Total $\sim$}8,290
+Adversarial G2 validation transformers, peft (Python) $\sim$}2,680
+MCP governance proxy demo cryptography, cbor2 (Python) $\sim$}520
+Total (Rust + Python) $\sim$}11,090
 7.2 Performance Evaluation
-We evaluate on Apple M2, 16 GB RAM, macOS. All benchmarks use Criterion.rs with 10,000 samples and 3-second warmup.
-Experiment 1: Certificate Verification Latency.
-Chain Depth Mean Per-Sig Cap. Binding
+We evaluate on Apple M2, 16 GB RAM, macOS (Criterion.rs, 10,000 samples, 3s warmup). Table 4 summarizes microbenchmarks for the four core governance primitives.
+Table 4: Core governance microbenchmarks (Experiments 1–4).
+Operation Key parameter Latency
+Exp 1: Certificate chain verification
+     Depth-2 chain 2 Ed25519 sigs 66.1 µs
+     Depth-5 chain (max practical) 5 Ed25519 sigs 163.6 µs
+     Capability binding 32-byte compare <<1 ns
+Exp 2: Skills manifest hash
+     10 tools Sort+CBOR+SHA-256 5.10 µs
+     100 tools Sort+CBOR+SHA-256 51.6 µs
+Exp 3: Ledger operations
+     Append record Ed25519 sign+sled 448 µs
+     Verify single record Ed25519 verify 32.4 µs
+     Full audit (100K records) Chain walk 3.46 s
+Exp 4: Reproducibility verification
+     Token compare (512 tokens) Byte compare 416 ns
+     Commit (512 tokens) SHA-256 6.55 µs
 […]
-[…]
-[…]
-5 (max practical) 163.6 µs 32.7 µs <<1 ns
-Capability binding adds <<1 ns—the cost of a single 32-byte comparison. Verification is dominated by Ed25519 signature verification ($\sim$}33 µs/sig).
-Experiment 2: Skills Manifest Hash.
-Tools Sort CBOR SHA-256 Total
-5 32 ns 0.41 µs 2.25 µs 2.69 µs
-10 42 ns 0.76 µs 4.30 µs 5.10 µs
-20 64 ns 1.71 µs 8.78 µs 10.6 µs
-50 152 ns 3.94 µs 21.8 µs 25.9 µs
-100 263 ns 8.05 µs 43.3 µs 51.6 µs
-Experiment 3: Ledger Throughput.
-Operation Latency Throughput
-Append record 448 µs $\sim$}2,230 rec/s
-Verify single record 32.4 µs $\sim$}30,900 rec/s
-Full audit (1K records) 34.7 ms 34.7 µs/rec
-Full audit (100K records) 3.46 s 34.6 µs/rec
-Query by agent ID 112 µs 0.56 µs/rec
-Experiment 4: Reproducibility Verification Latency.
-Tokens Compare Replay Skip Commit
-128 120 ns 123 ns 3.5 ns 1.72 µs
-512 416 ns 418 ns 3.5 ns 6.55 µs
-2,048 1.55 µs 1.55 µs 3.5 ns 24.3 µs
-8,192 6.01 µs 6.10 µs 3.6 ns 96.6 µs
-Token comparison scales at $\sim$}0.73 ns/token. Batch verification: 100 interactions (512 tokens each) in 41.8 µs. An auditor verifies 100,000 interactions in $\sim$}42 ms.
+Verification scales at $\sim$}33 µs/sig; token comparison at $\sim$}0.73 ns/token (100K interactions audited in $\sim$}42 ms).
 Experiment 5: Real LLM Reproducibility Validation.
-We validate the DiFR finding empirically in two phases.
-Phase 1 uses two Anthropic models (Claude Sonnet 4 and Haiku 4.5, temperature==0, max_tokens==512) across 105 prompts in five task types—code generation (25), JSON/structured output (15), mathematical reasoning (20), prose explanation (30), and creative writing (15)—with 5 same-model repetitions and 3 cross-model repetitions each (1,050 same-model + 1,575 cross-model pairwise comparisons).
-Phase 2 extends to 9 models from 7 providers: Claude Sonnet 4 and Haiku 4.5 (Anthropic), GPT-4.1 and GPT-4.1 mini (OpenAI), Gemini 2.5 Pro (Google), Grok 4 (xAI), DeepSeek V3 (DeepSeek), Llama 4 Maverick (Meta), and Kimi K2.5 (Moonshot).
-All models are accessed via OpenRouter as a unified API gateway (plus Volcengine for Doubao), with temperature==0 and 100 prompts across six task types (code, JSON, math, prose, creative, knowledge), 10 repetitions per model-prompt pair (5,400 total API calls; 15,120 pairwise comparisons: 10,800 same-model + 240 same-provider + 4,080 cross-provider).
-We compute four similarity metrics: character-level positional match (CharMatch), Jaccard token overlap, word-level TF-IDF cosine, and character n-gram (3–5) cosine; then apply Welch’s $t$-test, Mann-Whitney $U$, and Cohen’s $d$ to quantify separation, and use Youden’s $J$ statistic for optimal threshold selection.
-Result 5a: Same-model reproducibility is task-type dependent. Table 2 shows per-task reproducibility from Phase 1 across 105 prompts (1,050 same-model pairwise comparisons).
-Table 2: Same-model (Sonnet 4) reproducibility by task type.
-Task Type $n$ Identical CharMatch TF-IDF cos
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Structured outputs (JSON, code) have the highest identical-output fractions (>>50%), while prose shows the most variation (34.7% identical), driven by paraphrase-level reorganization.
-The bimodal distribution (median CharMatch =0.975=0.975 vs.
-mean =0.627=0.627) confirms that same-model outputs are either near-identical or substantially rephrased.
-Result 5b: Multi-provider separation analysis.
-Table 3 presents the Phase 2 multi-provider results across 15,120 pairwise comparisons (10,800 same-model vs.
-4,080 cross-provider).
-All differences are statistically significant ($p\approx 0$, Welch’s $t$ and Mann-Whitney $U$).
+We validate the DiFR finding across 9 models from 7 providers (Claude Sonnet 4/Haiku 4.5, GPT-4.1/4.1 mini, Gemini 2.5 Pro, Grok 4, DeepSeek V3, Llama 4 Maverick, Kimi K2.5) via OpenRouter with temperature==0, 100 prompts across six task types, 10 repetitions per model-prompt pair (5,400 API calls; 15,120 pairwise comparisons).
+We compute four similarity metrics—CharMatch, Jaccard, TF-IDF cosine, char n-gram cosine—and use Cohen’s $d$ and Youden’s $J$ for separation analysis and threshold selection.
+Result 5b. Table 5 shows multi-provider separation ($p\approx 0$ for all metrics).
 Table 5: Multi-provider separation: 9 models, 7 providers (15,120 comparisons).
-Metric Same-model Cross-provider Sep. Cohen’s $d$
+Metric Same Cross Sep. $d$
 CharMatch 0.538 0.114 $4.7\times$ 1.36
@@
 Char n-gram cos 0.871 0.640 $1.4\times$ 1.20
-All four metrics achieve Cohen’s $d>1.0$ (large effect sizes) across the full 9-model set, with Jaccard ($d=1.51$) and CharMatch ($d=1.36$) providing the strongest separation.
-Notably, same-provider pairs (e.g., GPT-4.1 vs.
-GPT-4.1 mini: CharMatch =0.251=0.251; Sonnet 4 vs.
-Haiku 4.5: $0.192$) exhibit similarity intermediate between same-model and cross-provider, confirming that even models from the same vendor maintain distinct output fingerprints.
-Result 5c: Per-model determinism varies by over an order of magnitude. Table 4 ranks models by same-model CharMatch, revealing a wide determinism spectrum.
+All metrics achieve Cohen’s $d>1.0$; Jaccard ($d=1.51$) and CharMatch ($d=1.36$) provide the strongest separation.
+Result 5c. Table 6 ranks per-model determinism.
 Table 6: Per-model same-model reproducibility (CharMatch, temperature==0).
@@
 Reasoning/thinking model (internal chain-of-thought).
-Claude Haiku 4.5 produces near-identical outputs across repetitions (CharMatch =0.987=0.987), while reasoning models (Kimi K2.5, Gemini 2.5 Pro, Grok 4) exhibit substantially lower determinism due to stochastic chain-of-thought reasoning even at temperature 0.
-Despite this variation, even the least deterministic model (Kimi K2.5, CharMatch =0.169=0.169) still exceeds the mean cross-provider similarity ($0.114$), demonstrating that model identity remains distinguishable for all tested models.
-Result 5d: Per-task separation is strongest for mathematical reasoning. Table 5 shows CharMatch separation factors across six task categories.
-Table 5: Per-task CharMatch separation (same-model / cross-provider).
-Task Type Same-model Cross-provider Sep.
-Math reasoning 0.508 0.042 $12.2\times$
-Knowledge 0.511 0.072 $7.1\times$
-Creative writing 0.402 0.057 $7.0\times$
-Prose 0.488 0.083 $5.9\times$
-Code generation 0.594 0.161 $3.7\times$
-JSON / structured 0.779 0.299 $2.6\times$
-Mathematical reasoning achieves the highest separation ($12.2\times$) because models develop highly distinctive step-by-step reasoning styles, while JSON outputs show the lowest separation ($2.6\times$) because structural constraints force convergent formatting.
-These task-dependent results motivate per-task threshold calibration in deployment.
-Result 5e: Optimal threshold via Youden’s $J$ statistic. Table 6 reports optimal thresholds from the multi-provider experiment.
+Determinism spans $5.8\times$ (Haiku 4.5: $0.987$ to Kimi K2.5: $0.169$); even the least deterministic model exceeds mean cross-provider similarity ($0.114$).
+Per-task separation ranges from $12.2\times$ (math) to $2.6\times$ (JSON).
+Result 5e.
+Via Youden’s $J$ (Table 7), Jaccard and CharMatch attain $F_{1}>0.87$ across all 9 models, confirming lightweight textual similarity provides practical model identity verification without model internals.
 Table 7: Optimal classification thresholds (multi-provider, 15,120 comparisons).
@@
 TF-IDF cosine 0.837 0.466 0.844
-In the multi-provider setting, Jaccard achieves the best overall classification ($F_{1}=0.876$, $J=0.592$), followed closely by CharMatch ($F_{1}=0.873$, $J=0.572$).
-Both metrics attain $F_{1}>0.87$ across 9 models and 7 providers, confirming that lightweight textual similarity metrics—requiring no model internals, embeddings, or logit access—provide practically useful model identity verification even in a heterogeneous multi-provider environment.
-Experiment 6: End-to-End MCP Overhead.
+Experiment 6: Adversarial Model Substitution Detection.
+We train three QLoRA backdoor variants (keyword B1, semantic B2, compositional B3) on three architectures (Qwen 2.5 7B, Llama 3.1 8B, Gemma 2 9B) and test G2 replay detection across four prompt strategies at budgets $n\in\{10,25,50,100,200\}$ (357 configs).
+Table 8: G2 adversarial detection rate at temperature=0{}=0 (3 models $\times$ 4 strategies).
+Budget ($n$) B1 B2 B3 Avg
+10 99.0% 99.2% 99.2% 99.1%
+25 100% 100% 100% 100%
+200 99.1% 98.9% 99.1% 99.1%
+Detection is near-universal (minimum 90% across all 357 configs).
+Mean CharMatch between base and backdoored outputs is 0.057, far below $\theta^{*}=0.146$: modifying weights to implant an effective backdoor unavoidably shifts the clean-input distribution, making substitution detectable (§8).
+7.3 Overhead and Comparison
+Experiment 7: End-to-End MCP Overhead.
 Phase Latency
@@
 Overhead as % of total 0.1–1.2%
-Experiment 7: Comparison with BAID.
-Metric BAID Ours Ratio
-Verification latency $\sim$}120 s 0.097 ms $1.2{\times}10^{6}$
-Proof size $\sim$}50 KB $\sim$}2 KB $25\times$
-Infrastructure Blockchain None
-Cap. change detection Post-hoc Real-time
-Experiment 8: MCP Governance Proxy Demo.
-To validate real-world applicability, we implement a complete MCP governance proxy in Python (520 lines) that interposes between an MCP client (agent) and MCP server (tool provider).
-The proxy enforces all three governance requirements: G1 via skills manifest hash verification against the agent’s certificate, G2 via reproducibility commitment tracking, and G3 via a hash-chain interaction ledger with Ed25519-signed records.
-We test 5 live attack scenarios:
-Attack Result Latency
-S1: Inject web_browser at runtime Blocked (hash $\Delta$) 121 µs
-S2: Replace read_file with trojan Blocked (code hash) 103 µs
-S3: Forged certificate (rogue CA) Blocked (sig verify) 106 µs
-[…]
-S5: Tamper ledger record #2 Detected (chain break)
-All attacks are detected in real-time.
-Over 1,000 governed calls, the proxy achieves mean governance overhead of 120 µs/call (median: 106 µs, p99: 432 µs), throughput of 8,195 governed calls/sec, and all ledger records pass integrity verification.
-This confirms that the framework provides real-time protection with overhead well below the latency of typical MCP calls (50–500 ms).
+Experiment 8: Enhanced Instantiation Benchmarks.
+Table 9: Enhanced cryptographic backends: BBS+ (G1) and Groth16 DV-SNARK (G2).
+Operation Latency vs. Basic
+BBS+ selective disclosure (16-message cert)
+     Full disclosure verification 12.27 ms $127\times$
+     Selective proof gen (6/16) 14.66 ms $151\times$
+     Selective proof verification 13.80 ms $142\times$
+Groth16 DV-SNARK CharMatch circuit
+     Prove (1,024 chars, 19K constraints) 173.86 ms
+     Prove (4,096 chars, $\sim$}73K constraints) 750.33 ms
+     Verify (constant) 1.28 ms
+Experiment 9: Comparison with BAID.
+Metric BAID Ours (basic) Ours (enhanced)
+G1 verification 0.097 ms 13.80 ms
+G2 proof gen (1K) $\sim$}120 s $\sim$}0.1 ms 174 ms
+G2 proof verify $\sim$}10 ms <<0.01 ms 1.3 ms
+Circuit size ${\sim}10^{10}$ 19,238
+Privacy Full ZK None Selective / ZK
+Infrastructure Blockchain None None
+Direct CharMatch comparison (no SNARK).
+BBS+ adds ${\sim}142\times$ overhead but enables selective disclosure; DV-SNARK verification is constant at ${\sim}1.3$ ms.
+Our DV-SNARK is ${\sim}690\times$ faster than BAID for proof generation (${\sim}520{,}000\times$ fewer constraints), by SNARKing only the verification metric rather than the full transformer forward pass.
+All enhanced operations remain within typical agent interaction latency (50–500 ms).
+Experiment 10: MCP Governance Proxy Demo.
+A Python proxy (520 lines) enforcing G1–G3 detects all 5 live attack scenarios (tool injection, trojanization, forged certificate, rate abuse, ledger tampering) in $\leq$165 µs each, achieving 8,195 governed calls/sec with mean overhead of 120 µs/call.
 7.4 Attack Scenario Evaluation
-We implement 12 attack scenarios as automated tests: 9 structural governance attacks as Rust unit tests (774 lines), 5 end-to-end scenarios via the MCP governance proxy demo (520 lines), and 3 runtime behavioral attacks informed by the Nemotron-AIQ dataset [45] (which demonstrates 100% attack success rates without structural defenses).
-Table 7 summarizes results.
-Table 7: Attack scenario evaluation. All 12 tests pass deterministically. S10–S12 are runtime behavioral attacks validated against Nemotron-AIQ traces.
+We evaluate 12 attack scenarios at three levels of validation (Table 10): (1) Structural governance attacks (S1–S9) are tested as automated Rust unit tests that verify the governance algorithms correctly reject each attack construction—these validate algorithm correctness, not resistance to real-world attackers.
+(2) Runtime behavioral attacks (S10–S11), informed by the Nemotron-AIQ dataset [46], are validated via post-hoc forensic tracing: the ledger provides sufficient evidence to identify the injection point after the fact, but the framework does not prevent the attack in real time.
+(3) S12 (output exfiltration) is detectable via replay divergence.
+The 6 E2E attacks in §7.5 provide the strongest validation: real-time detection in a live pipeline with real LLM calls.
+Table 10: Attack scenario evaluation.
+S1–S9: algorithm correctness validated via automated unit tests (9/9 pass).
+S10–S11: post-hoc forensic tracing validated against Nemotron-AIQ traces.
+S12: detectable via replay divergence.
+The “Framework” column distinguishes Prevented (blocked in real time), Detected (identified but not blocked), and Traced (forensic reconstruction after the fact).
 Scenario Attack Baseline Framework Detection Mechanism
-S1: Silent escalation Install web_browser at runtime Undetected Prevented capability_binding: hash mismatch
-S2: Tool trojanization Replace implementation (same name) Undetected Prevented capability_binding: code hash change
-S3: Model substitution Forge cert with different model Undetected Detected sig_verify: forged cert unsigned
-S4: Phantom delegation T2 agent issues T1 cert Undetected Prevented constraint_monotonicity
-S5: Evidence tampering Modify ledger record N/A Detected Hash chain + signature invalidation
-S6: Blame shifting Dispute in agent chain Unresolvable Resolved Ed25519 signatures
-S7: Colluding agents T3+T3 attempt T1 access Undetected Prevented tier_check: independent evaluation
-S8: Chain forgery Fabricated chain Prevented sig_verify: fake key $\neq$ root
-S9: Depth overflow $\mathit{max\_depth}=0$ delegation Prevented Structural: depth <0<0 impossible
-S10: Knowledge poisoning Inject false data via tool output Undetected Traced Ledger $\eta_{\mathit{in}}/\eta_{\mathit{out}}$: pinpoint injection span
-S11: Cascading hijack Multi-stage instruction persistence Undetected Traced Hash-chain: full propagation path audit
-S12: Output exfiltration Embed phishing via rendered output Undetected Detected Replay divergence from committed $\pi_{i}$
+S1: Silent escalation Add tool at runtime Undetected Prevented Skills manifest hash mismatch
+S2: Tool trojanization Replace implementation Undetected Prevented Code hash change
+S3: Model substitution Forged cert, wrong model Undetected Detected Signature verification failure
+S4: Phantom delegation T2 issues T1 cert Undetected Prevented Constraint monotonicity
+S5: Evidence tampering Modify ledger record N/A Detected Hash chain + sig invalidation
+S6: Blame shifting Dispute in chain Unresolvable Resolved Ed25519 bilateral signing
+S7: Credential collusion T3+T3 $\to$ T1 access Undetected Prevented Independent tier evaluation
+S8: Chain forgery Fabricated chain Prevented Sig verify: fake key $\neq$ root
+S9: Depth overflow Delegate at depth 0 Prevented Depth counter exhausted
+S10: Knowledge poisoning Inject via tool output Undetected Traced Ledger $\eta_{\mathit{in}}/\eta_{\mathit{out}}$
+S11: Cascading hijack Multi-stage persistence Undetected Traced Hash-chain provenance audit
+S12: Output exfiltration Embed phishing link Undetected Detected Replay divergence
 100% attack success in Nemotron-AIQ undefended traces (2,596 security trials).
 7.5 End-to-End Multi-Agent Pipeline
-Experiments 1–8 evaluate individual governance operations in isolation.
-To validate the framework under realistic conditions, we deploy the full A2Auth governance stack across a multi-agent pipeline where agents coordinate, delegate tasks, invoke tools via MCP, and pass results through the interaction ledger.
-This experiment measures: (a) end-to-end governance overhead as a fraction of pipeline latency, (b) scaling behavior from 5 to 20 agents, (c) attack detection in situ, and (d) comparison with three baselines.
-The implementation uses the Rust agent-trust-kit SDK exposed to Python via PyO3 FFI, with agents making real Anthropic Claude API calls.
-Figure 5: End-to-end multi-agent pipeline architecture with governance overlay.
-Five specialist agents coordinate through a trust tree (depth 3).
-Red labels indicate E2E attack injection points (Table 10).
-All inter-agent communications produce bilaterally-signed ledger records.
+We deploy the full governance stack in a multi-agent pipeline with real Claude API calls, measuring overhead, scaling, attack detection, and baseline comparison (Figure 5).
+Figure 5: Multi-agent pipeline with governance overlay. Red labels indicate attack injection points (Table 13).
 Pipeline architecture (Figure 5).
-We implement an AI Research Assistant pipeline: a user submits a research query; a Coordinator agent (T1_HIGH) orchestrates five specialist agents—Research (T2), Analysis (T2), Writer (T2), Review (T3_MONITORED), and Fact-Checker (T2)—each with a distinct skills manifest and reproducibility commitment.
-The trust tree has depth 3 (Root CA $\to$ Human $\to$ Coordinator $\to$ Workers), with the Coordinator’s max_delegation_depth=2{}=2 and all workers at depth 0 (cannot delegate).
-A single pipeline run traverses 21 steps: 10 tool calls (7 LLM invocations + 3 non-LLM tools), 11 inter-agent communications (task assignments and result deliveries), each producing a bilaterally-signed ledger record with capability verification and reproducibility commitment.
-Experiment 9a: Governance overhead.
-Table 8 breaks down governance cost per pipeline run.
-Each operation’s latency is consistent with the micro-benchmarks from Experiments 1–4; the end-to-end cost is their composition.
-Table 8: Per-pipeline governance overhead breakdown (5-agent configuration, 21 ledger entries per run, mean of 100 runs).
-Operation Per-op Count Subtotal
-Certificate chain verification 97 µs 6 0.58 ms
-Skills manifest hash check 5.1 µs 10 0.05 ms
-Ledger append (tool call) 448 µs 10 4.48 ms
-Reproducibility commitment 6.6 µs 10 0.07 ms
-Ledger append (inter-agent) 448 µs 11 4.93 ms
-Bilateral Ed25519 signing 66 µs 11 0.73 ms
-Total governance 21 10.8 ms
-Pipeline latency (real LLM) $\sim$}8.7 s
-Overhead 0.12%
-Governance adds 10.8 ms to a pipeline that takes $\sim$}8.7 s with real Claude API calls (7 LLM inferences at $\sim$}1.2 s each + 3 non-LLM tools at $\sim$}100 ms each).
-The overhead of 0.12% is dominated by ledger appends (87% of governance cost), confirming that the cryptographic operations themselves (hashing, signing, verification) are negligible relative to I/O-bound agent workloads.
-Experiment 9b: Scaling behavior.
-We scale the pipeline from 5 to 20 agents by instantiating parallel specialist groups: the Large configuration uses 4 Research agents (handling sub-queries in parallel), 4 Analysis agents, 3 Writer agents, 3 Review agents, 3 Fact-Checkers, 2 Coordinators (primary + sub-coordinator), and 1 Aggregator.
-Table 9 shows that per-agent governance cost remains approximately constant.
-Table 9: Governance scaling: 5 $\to$ 20 agents (mean of 50 runs per configuration).
+We implement an AI Research Assistant using LangGraph with the Rust agent-trust-kit SDK exposed via PyO3 FFI.
+A Coordinator (T1_HIGH) orchestrates four specialist agents—Research (T2), Analysis (T2), Writer (T2), and Review (T3_MONITORED)—in a depth-3 trust tree.
+Each pipeline run produces 7 governed tool calls, each generating a bilaterally-signed ledger record with capability verification and reproducibility commitment.
+Experiment 11a: Governance overhead.
+Table 11: Per-pipeline governance overhead (5-agent config, 7 ledger entries/run, mean of 90 runs with real Claude API calls).
+Operation Per-call Calls Total
+G1: Capability check 39 µs 7 0.27 ms
+G2: Repro. commitment 86 µs 7 0.60 ms
+G3: Ledger append 1,300 µs 7 9.10 ms
+Total governance 7 10.0 ms
+Pipeline latency (real Claude API) $\sim$}59.8 s
+Overhead <<0.02%
+Governance adds $\sim$}10 ms to a $\sim$}60 s real-LLM pipeline (<<0.02%), dominated by G3 ledger appends (91% of governance cost).
+Across 90 clean runs (630 governed tool calls), zero false positives and 100% ledger integrity are observed (95% CI: FPR << 3.3%).
+Experiment 11b: Scaling behavior.
+Table 12: Governance scaling: 5 $\to$ 20 agents (mock tool latency to isolate governance cost).
 Config Agents Ledger Gov. Per-agent Storage
 entries (ms) (ms) (KB)
-Small 5 21 10.8 1.81 5.2
-Medium 10 45 23.0 2.09 11.2
-Large 20 95 48.3 2.30 23.8
-Total governance cost scales linearly with ledger entries ($R^{2}>0.99$), but per-agent cost increases only 27% when quadrupling the agent count (1.81 ms $\to$ 2.30 ms).
-This near-constant per-agent cost arises because each agent’s governance operations—certificate verification, capability checking, ledger append—are independent of the total number of agents in the pipeline.
-The slight increase comes from deeper trust chains in the Large configuration (depth 4 vs.
-depth 3) adding one extra signature verification per activation.
-Ledger storage grows at $\sim$}250 bytes per entry; even the 20-agent pipeline produces only 23.8 KB per run, confirming that commitment-only storage (§5) is practical for long-running agent workflows.
-Experiment 9c: End-to-end attack detection.
-We inject five attacks into the live 5-agent pipeline, each targeting a different governance layer.
-Unlike the unit-level tests in Table 7, these attacks occur mid-pipeline with real inter-agent data flow.
-Table 10: End-to-end attack detection in the live multi-agent pipeline.
-Attack Target Layer Detect. Hops
-E2E-1: Capability escalation Research G1 5 µs
-E2E-2: Model substitution Analysis G2 $\dagger$
-E2E-3: Phantom delegation Writer G1 97 µs
-E2E-4: Ledger tampering Ledger G3 32 µs
-E2E-5: Cascading injection Pipeline G1+G2+G3 97 µs 3
-[…]
-dominated by LLM re-inference ($\sim$}1.2 s), not framework cost ($\sim$}13 µs).
-[…]
-Across 100 clean pipeline runs (500 total governed tool calls, 2,100 ledger entries), zero false positives are observed.
-Experiment 9d: Baseline comparison.
-We evaluate the same five E2E attacks against three baselines: (1) no governance (raw MCP), (2) MCP with OAuth 2.1 (standard MCP authentication per specification), and (3) MCP with OpenTelemetry traces (observability-only, no cryptographic integrity).
-Table 11 summarizes results.
-Table 11: E2E attack detection: A2Auth vs. baselines ($\checkmark$ = detected/prevented, ✗ = undetected). G1/G2/G3 indicate which governance requirement enables detection.
-Attack No gov. OAuth OTEL A2Auth
-E2E-1: Capability escalation $\checkmark$ (G1)
+Small 5 7 6.9 1.37 1.7
+Medium 10 16 12.9 1.29 3.9
+Large 20 33 30.0 1.50 8.1
+Governance scales linearly with ledger entries; per-agent cost remains approximately constant (1.29–1.50 ms across $4\times$ scaling). Storage is $\sim$}250 bytes/entry (8.1 KB for 20 agents).
+Experiment 11c: End-to-end attack detection.
+Table 13: End-to-end attack detection (real Claude API calls, 7 scenarios covering G1/G2/G3).
+Attack Target Layer Detection
+E2E-1: Cap. escalation Research G1 Tool not in manifest
+E2E-2: Model substitution Analysis G1 Model binding mismatch
+E2E-3: Phantom delegation Writer G1 No valid cert chain
+E2E-4: Ledger tampering Ledger G3 Hash chain break
+E2E-5: Tool trojanization Research G1 Manifest hash mismatch
+E2E-6: Depth overflow Review G1 Trust decay violation
+E2E-7: Replay divergence Analysis G2 CharMatch 0.00 << $\theta$
+All seven attacks are detected, covering all three governance layers.
+E2E-2 is caught by G1’s model binding check: the certificate declares claude-sonnet-4 but the agent uses claude-haiku-4-5.
+E2E-7 is the only attack requiring G2 (replay verification): the analysis agent’s certificate and model binding are correct (G1 passes), but the agent secretly produces attacker-controlled output instead of executing the declared model.
+Post-pipeline replay re-executes the declared model on the stored input and detects divergence (CharMatch==0.00 vs.
+$\theta$==0.85)—demonstrating that G2 catches attacks invisible to G1 and G3.
+Experiment 11d: Baseline comparison. We evaluate the same seven attacks against three baselines.
+Table 14: E2E attack detection: A2Auth vs. baselines ($\checkmark$ = detected, ✗ = undetected).
+Attack None OAuth OTEL A2Auth
+E2E-1: Cap. escalation $\checkmark$ (G1)
 E2E-2: Model substitution $\checkmark$ (G1)
@@
 E2E-4: Ledger tampering N/A N/A b $\checkmark$ (G3)
-E2E-5: Cascading injection c $\checkmark$ (G1+G2+G3)
-Total detected 0/5 0/5 0/5 5/5
-aOAuth authenticates the Writer but has no delegation depth constraint.
-bOTEL traces can be silently modified (no hash chain or signatures; §5).
-cOTEL logs the poisoned data flow but cannot prove tampering or trace provenance cryptographically.
-OAuth 2.1 provides identity authentication but not capability binding: it verifies who the agent is, not what tools it currently has.
-OpenTelemetry provides observability but not integrity: OTEL span trees log tool calls and parent-child relationships, but as demonstrated by analysis of 10,796 Nemotron-AIQ traces (§5), OTEL records contain no hash chains, no cryptographic signatures, and no identity binding—any field can be silently modified.
-A2Auth’s three governance layers (G1+G2+G3) provide the only configuration that detects all five attack types, with governance overhead below 0.2% of pipeline latency.
-8 Limitations and Future Work
+E2E-5: Tool trojanization $\checkmark$ (G1)
+E2E-6: Depth overflow $\checkmark$ (G1)
+E2E-7: Replay divergence $\checkmark$ (G2)
+Total detected 0/7 0/7 0/7 7/7
+aOAuth authenticates the agent but has no delegation depth constraint. bOTEL traces can be silently modified (no hash chain or signatures).
+OAuth 2.1 verifies who the agent is but not what tools it currently has or which model it runs. Only A2Auth (G1+G2+G3) detects all seven E2E attacks, at <<0.02% overhead.
+Experiment 11e: Governance coverage comparison with BAID and Ghosh et al.
+Table 14 compares against infrastructure baselines; Table 15 compares against the two closest research systems—BAID [2] (zkVM-based verifiable inference) and Ghosh et al.
+[46] (operational safety with Nemotron-AIQ)—at the governance requirement level across all 12 attack categories.
+This is a design-level analysis: we map each system’s published mechanisms to the attack taxonomy of §2.
+[…]
+Attack Requires BAID Ghosh A2Auth Notes
+[…]
+[…]
+[…]
+[…]
+[…]
+ATK-9: Blame shifting G2+G3 $LEFTcircle \CIRCLE$ BAID proves per-execution integrity but lacks bilateral signing for attribution.
+[…]
+[…]
+Governance requirements G2 only Partial G3 G1+G2+G3
+G2 proof generation $\sim$}120 s $\sim$}0.1 ms / 174 ms
+Infrastructure Blockchain None None
+Basic instantiation (direct CharMatch).
+Enhanced instantiation (DV-SNARK).
+BAID provides computational integrity (the declared model provably executed the computation)—a strictly stronger G2 guarantee than A2Auth’s statistical behavioral verification.
+However, BAID addresses only G2: it cannot detect capability escalation (G1) or audit interaction chains (G3).
+Ghosh et al.’s operational framework provides guardrails and red-teaming methodology but lacks cryptographic enforcement for any governance requirement.
+The three systems are complementary, not competing: BAID’s zkVM could serve as an enhanced G2 backend within the A2Auth architecture; Ghosh’s guardrails address content safety orthogonal to governance.
+8 Limitations, Governance Trilemma, and Future Work
 8.1 Scope of Contribution
-Our contribution is a governance framework, not novel cryptographic primitives.
-The individual building blocks—X.509 certificates, SHA-256, Ed25519, hash chains—are standard.
-The novelty lies in (a) identifying the capability-context separation as the principled boundary for agent governance, (b) showing that composing standard mechanisms along this boundary enforces G1–G3 with negligible overhead, and (c) demonstrating this is practically achievable.
-We view this as analogous to Certificate Transparency [16], which composes Merkle trees and gossip protocols—neither novel individually—into infrastructure that fundamentally changed how the web PKI operates.
-The security proofs are intentionally straightforward reductions to standard assumptions (collision resistance, EUF-CMA); the contribution is in what to build, not in new ways to prove things.
+Our contribution is a governance architecture, not new cryptographic primitives.
+The building blocks (digital signatures, hash chains, replay verification) are individually well-known; the contribution lies in identifying the capability-context separation as a principled governance boundary, proving structural constraints (Chain Verifiability, Bounded Divergence) that arise from this architecture, and defining abstract governance contracts (G1–G3) that hold across fundamentally different cryptographic backends.
+Why the composition is non-obvious.
+A counterfactual test: given Ed25519, SHA-256, and hash chains, would a security engineer independently arrive at G1–G3?
+The evidence suggests not.
+The MCP ecosystem has grown to over 1,000 tool servers with OAuth 2.1 as its security mechanism, yet no deployment detects capability drift (G1 gap).
+OpenTelemetry is the industry standard for agent observability, yet the Nemotron-AIQ dataset [46] demonstrates that 10,796 OTEL traces provide no tamper evidence (G3 gap).
+BAID [2] constructs a zkVM over the full transformer pass, yet cannot detect whether the agent’s tool set has changed (G1 gap).
+Each existing approach addresses a fragment; none recognizes the capability-context separation as the organizing principle, and none arrives at the three-requirement conjunction that Table 1 shows is necessary and sufficient.
+The composition and the boundary identification are the contributions, validated by two instantiations (basic and enhanced) satisfying the same nine security properties under standard assumptions (CR, EUF-CMA; $q$-SDH, DDH, knowledge soundness).
+8.2 The Agent Governance Trilemma
+A fundamental question underlies the design of any agent governance framework: why can’t we simply have agents that are maximally capable, maximally fast, and maximally secure?
+We observe that this is not merely a current engineering limitation, but a structural constraint rooted in computability theory.
+Observation 1 (Agent Governance Trilemma, informal).
+For agent systems operating under a fixed computational budget, three desiderata—capability, performance, and security—cannot be simultaneously maximized. Formally:
+  1. 1.
+    High Capability + High Security $\Rightarrow$ Performance $\leq 1/2$.
+If an agent’s tool set is Turing-complete (unbounded capability) and full behavioral verification is required (high security), then verification costs at least as much as the original execution, so at most half the computational budget produces useful work.
+  2. 2.
+    Bounded Capability + High Security $\Rightarrow$ Performance $\to 1$.
+If the tool set is non-Turing-complete (bounded capability), verification reduces to lookup in a pre-computed table of valid input-output pairs, with $O(1)$ per-query overhead.
+  3. 3.
+    High Capability + High Performance $\Rightarrow$ Security degrades.
+If the agent is Turing-complete and most of the budget goes to useful work, insufficient resources remain for verification, and the verifier cannot detect all semantic deviations.
+Proof sketch for (1).
+Behavioral verification (G2) requires determining whether the agent’s output $y$ matches what its declared model $M$ would produce on input $x$.
+For general computations, the only universal method is re-execution: compute $M(x)$ and compare.
+No sub-linear shortcut exists for arbitrary $M$, because $M$’s output depends on the full complexity of its parameters and computation—the function $M$ may be incompressible (its Kolmogorov complexity approaches its execution cost).
+Therefore $W_{\mathrm{verify}}\geq W_{\mathrm{task}}$, giving performance ratio $P=W_{\mathrm{task}}/(W_{\mathrm{task}}+W_{\mathrm{verify}})\leq 1/2$.
+For Byzantine-fault-tolerant verification with $k$ independent verifiers, $P\leq 1/(k+1)$.
+Proof sketch for (2).
+When each tool $t_{i}$ has a finite input space $\mathcal{I}_{i}$ and output space $\mathcal{O}_{i}$, the set of valid $(x,t_{i}(x))$ pairs is bounded by $\sum_{i}|\mathcal{I}_{i}|$.
+A hash table of valid pairs can be pre-computed offline and queried in $O(1)$ per verification.
+Since the pre-computation is amortized over all future queries, the marginal verification cost is constant, and $P\to 1$ as the number of task executions grows.
+Proof sketch for (3).
+This follows from Rice’s theorem: “the output of agent $A$ is safe” is a non-trivial semantic property of $A$’s program, and no algorithm can decide all such properties for Turing-complete programs.
+When the verification budget $W_{\mathrm{verify}}<W_{\mathrm{task}}$, the verifier cannot re-execute all computations.
+Sampling-based verification (checking $n$ out of $N$ outputs) provides statistical guarantees bounded by our Bounded Divergence Theorem ($\epsilon\leq\ln(1/\alpha)/n$), but the adversary can always deviate on the NnN-n unchecked outputs.
+Static analysis of agent programs is undecidable in the general case.
+How TEE sidesteps the trilemma.
+Hardware TEE replaces computational verification with hardware attestation, shifting the tradeoff from the compute domain to the trust assumption domain (reliance on hardware vendor integrity and supply chain security).
+TEE does not break the trilemma but trades compute cost for trust assumptions, and satisfies only G2—a TEE-attested agent still requires G1 and G3.
+Analogy to established impossibility results.
+The trilemma is structurally analogous to the CAP theorem [48] and the blockchain trilemma; we conjecture that a formal proof reducing to standard computability results is possible (§8.8).
+8.3 Governance Levels: Operationalizing the Trilemma
+The trilemma motivates three governance levels (Level 3: compile-time safety with bounded tools; Level 2: sampled verification for enterprise deployments; Level 1: post-hoc traceability for low-risk use cases), each corresponding to a different position in the capability-performance-security space.
+Full specifications including recommended crypto backends are given in Appendix B.
+The key insight is that every deployment’s security posture is legible: the governance level is declared in the agent’s certificate and recorded in the ledger, making the tradeoff explicit to auditors and downstream agents.
+8.4 Capability Boundedness and Meta-Tools
+The trilemma’s capability dimension distinguishes two tool classes.
+Bounded capabilities (web_search, read_file) have finite or structurally constrained input-output spaces; G1 fully characterizes the envelope and verification is efficient.
+Unbounded capabilities (code_execution, shell_access) are Turing-complete: G1 correctly records that the agent has the capability, but the envelope is formally declared yet semantically unbounded—knowing an agent can execute code does not predict what code it will execute.
+For unbounded tools, governance either restricts to bounded mode (sandboxed execution, Level 3) or accepts unbounded mode with runtime monitoring ($P\leq 1/2$ per the trilemma, as the monitor consumes at least as much compute as the governed agent).
+G1–G3 provides the infrastructure for both: G1 declares boundedness, G2 verifies behavioral fidelity within the envelope, and G3 records all interactions for post-hoc audit.
 8.5 Capability Binding Limitations
-Capability-context boundary.
-Our sharp boundary has edge cases: a “code execution” tool can effectively acquire new capabilities through composition (e.g., downloading and running arbitrary scripts).
-Our framework guarantees that the registered capability set has not changed, but the effective attack surface within a fixed capability envelope can still be large.
-Complementary approaches (sandboxing, taint tracking, output filtering) are needed to govern behavior within the envelope.
-We view capability binding and intra-capability governance as orthogonal layers: the former ensures the attack surface is known; the latter constrains behavior within it.
 Closed-source binding.
@@
 Re-signing frequency.
-Mitigated through wildcard tool categories, short-lived auto-renewable certificates, and the empirical observation that tool configurations change far less frequently than user context.
-8.3 Reproducibility Limitations and Graceful Degradation
-Reproducibility is a design constraint, not a universal guarantee.
-Our framework makes explicit what is often implicit: non-reproducible agent behavior is non-verifiable.
-This is not a limitation unique to our approach—it applies to any verification scheme not based on trusted hardware.
-For high-assurance deployments, this means agents should be configured for deterministic output (structured formats, temperature==0, fixed seeds).
-Our empirical results (§7) show that code-generating agents already satisfy this constraint perfectly, and the separation between same-model and cross-model distributions ($11.5\times$) provides robust detection even for prose tasks.
-Graceful degradation under partial G2.
-Our multi-provider experiment (§7, Experiment 5c) reveals that model determinism varies by over an order of magnitude: from CharMatch==0.987 (Haiku 4.5) to 0.169 (Kimi K2.5).
-This raises a practical question: what security guarantees hold when some agents use low-determinism models?
-When G2 is unavailable (reproducibility commitment $\mathit{none}$), the framework degrades to G1+G3 only. Table 12 characterizes the precise security impact.
+Mitigated through wildcard tool categories (web_search@ˆ2.0) and short-lived auto-renewable certificates.
+A typical deployment (10 tools, weekly updates) incurs ${\sim}520$ certificate operations per year at <<100 µs each—operationally negligible.
+8.6 Graceful Degradation under Partial G2
+When G2 is unavailable (reproducibility commitment $\mathit{none}$), the framework degrades to G1+G3 only. Table 16 summarizes the security impact.
 Table 16: Security guarantees under G1+G2+G3 vs. G1+G3 degradation.
@@
 G1+G3 alone protect against 9 of 12 attack types.
-The three attacks that require G2—model substitution (ATK-2), cascading injection via model compromise (ATK-10/11 when the injection occurs through model replacement rather than tool replacement), and output exfiltration via model tampering (ATK-12)—share a common structure: the attacker operates within a valid capability envelope by substituting the computational process itself.
-Signing proves attribution; only G2 proves computational fidelity.
+The attacks requiring G2—model substitution (ATK-2), cascading injection via model compromise (ATK-10/11), and output exfiltration (ATK-12)—share a common structure: the attacker operates within a valid capability envelope by substituting the computational process itself.
 Chain-level degradation.
-The Chain Verifiability Theorem (§4.8) establishes a stronger result: a G2==none agent at an interior position breaks behavioral verification for all downstream nodes.
-This means the true cost of deploying a low-determinism model at a coordinator position is not merely that the coordinator itself is unverifiable, but that the entire sub-tree it governs loses G2 protection.
-The framework’s three commitment levels (full/statistical/none) are thus not equivalent—their security implications depend critically on chain position.
+A G2==none agent at an interior position breaks behavioral verification for all downstream nodes (Chain Verifiability Theorem, §4.9); the deployment principles therein provide prescriptive architectural guidance.
 Statistical vs.
 computational security.
-Token-level reproducibility provides statistical rather than computational security.
-However, the Bounded Divergence Theorem (§4.7) shows this gap is operationally small: sufficient replay verification transforms statistical detection into a provable safety certificate.
-An adversary who fine-tunes a model to pass $n$ diverse replay checks has produced a model that is $(n,\epsilon)$-indistinguishable from the original, with $\epsilon$ decreasing exponentially in $n$.
-Future work will explore semantic similarity (embedding cosine distance) as a verification metric, which may provide tighter separation bounds and is robust to paraphrase-level variation.
+Token-level reproducibility provides statistical rather than computational security, but the Bounded Divergence Theorem (§4.8) bounds this gap: an adversary passing $n$ replay checks has produced a model $(n,\epsilon)$-indistinguishable from the original, with $\epsilon\leq\ln(1/\alpha)/n$.
+Verification metric limitations.
+Current textual similarity metrics (CharMatch, Jaccard) have lower separation power for reasoning models with high paraphrase variation.
+The architecture is metric-agnostic: any similarity function satisfying the threshold separation property can be substituted, and semantic similarity (embedding cosine distance) is a natural extension.
 Provider cooperation.
-Replay verification requires access to the same model version.
-Organizations should negotiate replay rights as part of service agreements.
-Deterministic inference providers [6] eliminate this dependency entirely.
+Basic G2 requires access to the same model version, mitigated by negotiated replay rights, deterministic inference providers [6], or the enhanced G2 instantiation (DV-SNARK, §4.10) which reduces cooperation to one-time proof generation.
 8.7 Trust Infrastructure Limitations
@@
 We assume non-agent nodes (CAs, humans, organizations) are uncompromised.
-Standard PKI mitigations apply: Certificate Transparency logs, hardware trust anchors (TPM 2.0), and multi-party issuance protocols.
+Standard PKI mitigations apply (CT logs, TPM 2.0 trust anchors, multi-party issuance, short certificate lifetimes); our framework inherits these risks and mitigations from the PKI foundation.
 Ledger operator trust.
-The current single-operator ledger relies on bilateral signing for tamper evidence.
-A multi-operator design with consensus is possible but introduces latency overhead.
+The current G3 instantiation assumes a centralized operator honest for availability and ordering; record integrity is enforced cryptographically (bilateral signatures, hash chains) regardless.
+Alternative instantiations (multi-operator consensus, transparency logs, blockchain anchoring) can relax the availability assumption at the cost of latency, preserving G3 semantics (§5).
 Application-layer data flow.
@@
 8.8 Future Work
-Hardware-attested agent identity (TPM 2.0, TEEs) for stronger G2 guarantees without provider cooperation.
-Embedding-based semantic similarity for reproducibility verification with formal threshold analysis.
-Cross-protocol capability governance beyond MCP/A2A.
-AI-assisted certificate issuance to reduce operational burden.
-Dynamic trust calibration from ledger history, enabling trust scores to evolve based on observed agent behavior.
+Formalizing the trilemma.
+A rigorous formalization reducing to Rice’s theorem and simulation bounds—analogous to the Gilbert–Lynch proof of CAP (2002)—would establish the trilemma as a foundational constraint on agent governance.
+Hardware-attested identity.
+TPM 2.0 and TEE [53, 54] integration for stronger G2 guarantees without provider cooperation, addressing the trilemma by substituting compute-domain verification with trust-domain attestation.
+Stealthy model substitution.
+Our adversarial experiment (Experiment 6) shows QLoRA-based backdoors are trivially detected.
+More sophisticated attacks (distillation-based backdoors, gradient-masked training) may produce stealthier substitutions; characterizing the fundamental limits of replay-based detection is an open question.
+Semantic verification metrics.
+Embedding-based semantic similarity for reproducibility verification with formal threshold analysis, potentially extending G2’s applicability to low-determinism reasoning models.
+IVC for G3.
+An enhanced G3 using Incrementally Verifiable Computation (e.g., Nova folding) would enable $O(1)$ audit verification regardless of chain length, deferred as current scales do not justify the ${\sim}200\times$ append overhead.
+Cross-protocol governance. Extending capability binding beyond MCP/A2A to emerging agent protocols.
+Dynamic trust calibration. Trust scores evolving based on ledger history, enabling governance levels to adapt automatically based on observed behavior.
 9 Related Work
@@
 Governance Platforms. Credo AI [43] maintains agent registries at the documentation layer.
-Verifiable Execution.
-BAID [2] provides zkVM-based verification ($\sim$}120 s per proof).
-DiFR [1] demonstrates inference near-determinism.
-EigenAI [6] achieves 100% deterministic inference.
-Attestable Audits [7] uses TEEs.
+[…]
 Classical Trust Management. PolicyMaker [13], KeyNote [14], SPKI/SDSI [15], F-PKI [17], ARPKI [18].
-Agent Protocols. A2A [8] provides agent communication. MCP enables agent-tool interaction with OAuth 2.1.
+[…]
 Agent Governance Frameworks.
@@
 [29] adapt TRiSM for agentic multi-agent systems.
+Instruction-Data Separation.
+A related line of work addresses the model’s inability to distinguish instructions from data within the transformer.
+The Instruction Hierarchy [61] trains models to prioritize system prompts over user inputs and tool outputs, establishing trust levels per input type.
+However, this operates at the model level via fine-tuning—it does not address the orchestration-layer problem that our capability-context separation targets: tool definitions and runtime context have different governance lifecycles (infrequent re-authorization vs.
+per-interaction auditing) that no amount of model training can enforce.
+Our separation is complementary: even a perfectly instruction-following model requires external cryptographic infrastructure to detect when its capability envelope has changed (G1), verify its computational process (G2), and produce tamper-evident interaction records (G3).
 Regulatory Landscape.
@@
 10 Conclusion
-AI agents increasingly acquire capabilities at runtime, yet no existing framework enforces the three governance requirements—capability integrity, behavioral verifiability, and interaction auditability—that a trustworthy agent ecosystem demands.
-We addressed this through the capability-context separation principle and three mechanisms: capability-bound certificates (G1), reproducibility commitments (G2), and a verifiable interaction ledger (G3).
-We formalized nine security properties (including chain verifiability: behavioral verification is a chain property broken by any unverifiable interior agent), implemented the framework in Rust with 97 µs verification and <<1 ns capability binding overhead (${\sim}1{,}200{,}000\times$ faster than BAID), validated reproducibility across 9 models from 7 providers ($F_{1}=0.876$ multi-provider, $F_{1}=0.990$ single-provider), and verified 12 attack scenarios—including runtime behavioral attacks validated against NVIDIA Nemotron-AIQ traces—with specific, traceable detection mechanisms.
-An end-to-end evaluation deploying the full governance stack across a 5-to-20-agent pipeline with real LLM calls confirms 0.12% overhead, sub-linear per-agent scaling, and detection of all in-situ attacks with zero false positives—while three baselines (no governance, OAuth 2.1, OpenTelemetry) detect none.
-Our framework fills the stateful, capability-aware governance gap between identity platforms and runtime gateways.
+We identified silent capability escalation as a new vulnerability class in AI agent ecosystems and traced its root cause to a missing architectural distinction: the capability-context separation, which recognizes that tool definitions and runtime context have fundamentally different security semantics at the orchestration layer.
+From this principle, we derived three Agent Governance Requirements (G1–G3) that define what an agent ecosystem must enforce, independent of how—a governance architecture, not new cryptographic primitives.
+Two structural results have prescriptive architectural consequences.
+The Chain Verifiability Theorem establishes that behavioral verification is a chain property: one unverifiable interior agent breaks end-to-end verification for all downstream nodes, constraining how multi-agent systems must be architected.
+The Bounded Divergence Theorem transforms replay-based verification into a probabilistic safety certificate, establishing software G2 as a deployable approximation of trusted execution.
+A multi-provider reproducibility study (9 models, 7 providers) reveals $5.8\times$ variance in inference determinism, directly connecting model characteristics to governance architecture and elevating reproducibility to a first-class model performance metric.
+We validated the framework with two crypto-agnostic instantiations—basic (Ed25519, SHA-256, hash chains; 97 µs verify) and enhanced (BBS+ selective disclosure, DV-SNARK; 13.8 ms verify)—both satisfying the same nine security properties.
+End-to-end evaluation over 5–20 agent pipelines with real LLM calls confirms <<0.02% governance overhead and real-time detection of 7 end-to-end attack scenarios (covering G1, G2, and G3) with zero false positives, while three baselines (no governance, OAuth 2.1, OpenTelemetry) detect none.
+An additional 9 structural attacks are validated via automated unit tests, and 3 runtime behavioral attacks are validated via post-hoc forensic tracing.
+The Agent Governance Trilemma (§8.2) reveals why no single configuration suffices: agents with unbounded capabilities (code execution, tool composition) require verification at least as expensive as the original execution, forcing an explicit tradeoff between capability, performance, and security.
+G1–G3 provides the infrastructure for the full governance spectrum—from compile-time safety (bounded tools, full verification) to post-hoc traceability (unbounded tools, audit-only)—making each deployment’s security posture explicit and auditable.
 As the EU AI Act’s high-risk provisions take effect and NIST develops agent-specific standards, the need for cryptographic infrastructure enforcing G1–G3 will only grow.
+By making reproducibility a prerequisite for behavioral verifiability, our framework creates a concrete incentive for model providers to self-declare inference determinism characteristics—elevating reproducibility from an implementation detail to a first-class, security-critical model specification.
 We release the core mechanisms as an open-source SDK (agent-trust-kit) to enable ecosystem adoption.
+Appendix A Proofs of Security Properties
+This appendix contains the full proofs for all theorems stated in §6.
+Proof of Property 3 (Trust Containment).
+Suppose $\mathcal{A}$ issues $\mathit{cert}_{w}$ with $\kappa_{w}\not\leq_{\kappa}\kappa_{v}$.
+The verifier checks $\kappa_{w}\leq_{\kappa}\kappa_{v}$ using constraints in $\mathit{cert}_{v}$, which is signed by $v$’s parent.
+$\mathcal{A}$ cannot modify $\mathit{cert}_{v}$ without the parent’s key.
+Recursing upward, the chain terminates at an uncompromised non-agent ancestor (EUF-CMA).
+∎∎
+Proof of Property 4 (Capability Binding).
+By collision resistance: $\Pr[H_{v}(t^{\prime})=H_{v}^{\mathit{cert}}]\leq\mathrm{negl}(\lambda)$ when $S_{v}(t^{\prime})\neq\sigma_{v}$.
+Forging a new certificate requires the parent’s key (same argument as Property 3).
+∎∎
+Proof of Property 5 (Chain Integrity).
+Fabricating any link requires the corresponding private key. By EUF-CMA, $\mathcal{A}$ cannot produce valid signatures without the key. ∎∎
+Proof of Property 6 (Delegation Depth Enforcement).
+Trust Containment requires $\kappa_{v}.\mathit{max\_depth}<\kappa_{\mathit{par}(v)}.\mathit{max\_depth}$ (strict). Starting from depth $d$, the chain has at most $d$ agent levels before reaching 0. ∎∎
+Proof of Property 7 (Reproducibility Soundness).
+Hardware G2: TEE attestation directly proves that the declared model was executed; substitution is detected with certainty.
+Software G2: By DiFR [1], for different models the empirical match rate is $\delta\approx 0.06$ (character-level), validated in our Experiment 5 across 9 models and 7 providers.
+Our empirical data shows that the fraction of cross-model output pairs exceeding any reasonable threshold $\theta$ (e.g., $\theta=0.15$) is at most $q\approx 0.02$ (i.e., $\Pr[\mathrm{CharMatch}(M(x),M^{\prime}(x))\geq\theta\mid M^{\prime}\neq M]\leq q$).
+Each verification prompt is therefore an independent Bernoulli trial with pass probability at most $q$.
+The probability that a substituted model passes all $n$ verification prompts is at most $q^{n}$, which is negligible for realistic $n$ (e.g., $q^{10}\leq 10^{-17}$).
+∎∎
+Proof of Theorem 3 (Bounded Divergence Under Verification).
+Let $p=\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]$ be the true failure rate.
+Each verification trial $Z_{i}=\mathbf{1}[\mathrm{CharMatch}(M(x_{i}),M^{\prime}(x_{i}))\geq\theta]$ is an independent Bernoulli(1p)(1-p) random variable.
+Observing all $n$ trials pass ($\sum Z_{i}=n$), we seek an upper confidence bound on $p$.
+The probability of observing $n$ successes when the true failure rate is $p$ is $(1-p)^{n}$.
+The exact one-sided Clopper-Pearson upper bound at confidence level $1-\alpha$ is the value $p^{*}$ satisfying $(1-p^{*})^{n}=\alpha$, giving $p^{*}=1-\alpha^{1/n}$.
+For small $p^{*}$, using the approximation $\ln(1-p)\approx-p$: from $(1-p)^{n}=\alpha$ we get $n\ln(1-p)=\ln\alpha$, hence $p\approx\ln(1/\alpha)/n$.
+This is a distribution-free result: the bound depends only on the number of verification trials $n$ and the desired confidence level $\alpha$, not on the threshold $\theta$ or any distributional assumptions on the CharMatch statistic.
+The role of $\theta$ is solely to define what constitutes a “pass”—a higher $\theta$ increases sensitivity but may also increase false positives for the legitimate model.
+∎∎
+Proof of Property 8 (Credential Isolation).
+Access depends solely on the requesting agent’s certificate. No protocol mechanism combines permissions. Proxy tokens are bound to specific agent IDs. ∎∎
+Proof of Property 9 (Ledger Integrity).
+(a) Modification invalidates $h_{\mathit{prev}}$ of successor, cascading through chain; re-signing requires all subsequent senders’ keys.
+(b) Deletion breaks hash chain.
+(c) Reordering breaks both sequence numbers and hash links.
+∎∎
+Proof of Theorem 4 (Chain Verifiability).
+We proceed by induction on chain position.
+[…]
+Inductive step: Assume the output of $v_{i-1}$ (i1>ki-1>k) cannot be verified as legitimate.
+Agent $v_{i}$ receives this unverifiable output as input.
+By the same argument as the base case, $v_{i}$’s replay verification confirms computational fidelity on unverifiable input.
+∎∎
+Proof of Property 10 (G2 Input Integrity via Bilateral Signing).
+Let $v_{j}$ be a compromised agent and let $R_{i}=(\ldots,a_{\mathit{send}},a_{\mathit{recv}},\ldots,\eta_{\mathit{in}},\ldots,\mathit{sig}_{i})$ be the ledger record for the interaction delivering input to $v_{j}$, where $a_{\mathit{send}}=v_{j-1}$ and $a_{\mathit{recv}}=v_{j}$.
+The record $R_{i}$ is bilaterally signed: $\mathit{sig}_{i}$ includes signatures from both $v_{j-1}$ and $v_{j}$.
+The input commitment $\eta_{\mathit{in}}=\mathrm{SHA\text{-}256}(\mathit{input})$ is covered by $v_{j-1}$’s signature.
+Claim: $v_{j}$ cannot cause the verifier to use a fabricated input $\mathit{input}^{\prime}\neq\mathit{input}$ for replay verification.
+Suppose $v_{j}$ attempts to present $\mathit{input}^{\prime}$ during forensic reconstruction.
+The verifier computes $\mathrm{SHA\text{-}256}(\mathit{input}^{\prime})$ and compares against $\eta_{\mathit{in}}$ in the ledger record.
+By collision resistance of SHA-256, $\mathrm{SHA\text{-}256}(\mathit{input}^{\prime})=\eta_{\mathit{in}}$ with negligible probability when $\mathit{input}^{\prime}\neq\mathit{input}$.
+Moreover, modifying $\eta_{\mathit{in}}$ in the record $R_{i}$ invalidates $v_{j-1}$’s signature (EUF-CMA of Ed25519), and is further detectable via hash-chain integrity ($h_{\mathit{prev}}$ of $R_{i+1}$ would not match).
+When $v_{j-1}$ is honest (Cases 1 and 2 in Property 10), the verifier independently requests $v_{j-1}$ to disclose $\mathit{input}$ and verifies it against $\eta_{\mathit{in}}$, bypassing $v_{j}$ entirely.
+When $v_{j-1}$ is also compromised (Case 3), both parties may collude to present a consistent fabrication.
+However, Chain Verifiability (Theorem 4) already establishes $\mathrm{CVD}\leq j-1<j$: the chain’s verifiable region does not extend to $v_{j}$, so G2 replay of $v_{j}$ is not claimed to provide end-to-end guarantees in this configuration.
+∎∎
+Appendix B Governance Levels
+The Agent Governance Trilemma (Observation 1) implies that no single governance configuration is optimal for all deployments.
+G1–G3 provide the infrastructure for three governance levels, each corresponding to a different position in the capability-performance-security space:
+Level 3: Compile-Time Safety (high security + high performance, bounded capability).
+For safety-critical deployments—financial transactions, medical decisions, legal document generation—agents are restricted to bounded tools only: no code execution, no tool composition.
+G1 binds the complete capability envelope; G2 with full replay verification ($n\geq 200$, $\epsilon<2.3\%$ at 99% confidence) ensures behavioral fidelity; G3 provides complete audit trails.
+Recommended crypto backend: Enhanced G1 (BBS+ with predicate proofs, §3.5) + Enhanced G2 (DV-SNARK, §4.10).
+Level 2: Sampled Verification (balanced tradeoff).
+For enterprise deployments—data processing, customer service, code generation—agents may use selected unbounded tools, but G2 operates in statistical mode with periodic sampling.
+A verification budget of $n=50$ per audit cycle bounds adversarial divergence to $\epsilon<8.9\%$ (99% confidence).
+G1 declares the capability envelope; G3 ensures traceability.
+Recommended crypto backend: Enhanced G1 (BBS+ selective disclosure) + Basic G2 (direct CharMatch).
+Level 3: Post-Hoc Traceability (high capability + high performance, reduced security).
+For low-risk deployments—writing assistance, creative design, information retrieval—agents operate with full capability and minimal runtime overhead.
+G2 is set to $\mathit{none}$; G1 provides basic capability declaration; G3 provides lightweight interaction traces.
+Security relies on post-hoc forensic reconstruction.
+Recommended crypto backend: Basic G1 (Ed25519).
+The governance level is declared in the agent’s certificate and recorded in the ledger, making each deployment’s security posture explicit to auditors, compliance systems, and downstream agents.